### 24 Mar 2026 v1

# Spectral Structure of the Mixed Hessian of the Dispersionless Toda $\tau$-Function

Abstract.
We analyze the mixed Hessian of the dispersionless Toda $\tau$-function for the $s$-fold symmetric one-harmonic polynomial conformal map. The inverse branch exhibits two distinct thresholds: an analytic threshold $\zeta_{c}$, where the dominant square-root singularity reaches the circle of convergence, and a later geometric threshold $\zeta_{\mathrm{univ}}>\zeta_{c}$, where the map ceases to be univalent. We prove that the first spectral instability occurs already at $\zeta_{c}$. In each symmetry sector, the weighted subcritical realization has exactly one logarithmically diverging eigenvalue, whereas the remaining spectrum stays bounded and, after removal of the singular direction, converges to that of a compact limiting remainder. We further continue the corresponding scalar Gram functions beyond $\zeta_{c}$, showing that they admit a generalized hypergeometric description, a Cauchy–Stieltjes representation, and, for $1\leq p\leq s$, a realization as Weyl functions of bounded Jacobi operators. In particular, these scalar quantities remain finite at $\zeta_{\mathrm{univ}}$. This identifies analytic criticality, rather than loss of univalence, as the first spectral threshold of the Toda Hessian.
1. Introduction
Planar domains, conformal maps, and harmonic moments play a central role in several closely related subjects, including Laplacian growth, dispersionless integrable hierarchies, Green-function and Bergman-kernel theory, and planar matrix models [17, 24, 16, 13, 9, 25, 23]. A common feature of these theories is that the geometry of a simply connected domain may be encoded by its conformal map and by the harmonic moments of its complement. This makes it natural to ask not only how the domain itself degenerates near criticality, but also how the associated second-order response objects behave. The present paper studies one such object: the mixed Hessian of the dispersionless Toda free energy, equivalently the coefficient matrix of a mixed logarithmic kernel attached to the inverse conformal map.
Let $D\subset\mathbb{C}$ be a bounded simply connected domain with rectifiable boundary, and let
$f:\{|w|>1\}\to\mathbb{C}\setminus D,\qquad f(w)=rw+a_{0}+a_{1}w^{-1}+\cdots,$
be the unique exterior conformal map normalized by $f(w)/w\to r$ as $w\to\infty$. The parameter $r>0$ is the conformal radius of $D$, while the remaining coefficients encode the shape of the boundary. Let $w=w(z)$ denote the inverse branch near infinity, normalized by $w(z)=z/r+O(z^{-1})$.
Two bilinear kernels are naturally attached to $f$. The first is the holomorphic Faber–Grunsky kernel $\log\frac{f(w)-f(w^{\prime})}{r(w-w^{\prime})}$ whose coefficients control univalence through the classical Grunsky inequalities [5, 19]. By contrast, the present paper is concerned with the mixed holomorphic–antiholomorphic logarithmic kernel
(1.1) $\log\!\left(1-\frac{1}{w(z)\,\overline{w(z^{\prime})}}\right)=-\sum_{m,n\geq 1}H_{mn}\,\frac{(z/r)^{-m}}{m}\frac{(\bar{z}^{\prime}/r)^{-n}}{n},$
valid for $|z|$ and $|z^{\prime}|$ sufficiently large. It is obtained by pulling back the canonical mixed logarithmic kernel of the exterior disk under the inverse conformal map. The coefficients $H_{mn}$ are the main object of the present paper.
In the geometric realization of the dispersionless 22D Toda hierarchy, the exterior conformal map evolves under commuting flows whose times $(t_{0},t_{1},t_{2},\dots;\bar{t}_{1},\bar{t}_{2},\dots)$ are identified with the harmonic moments of $\mathbb{C}\setminus D$. In a standard normalization, $t_{0}=(1/\pi)\int_{D}d^{2}z$ is the area of $D$ divided by $\pi$, while for $n\geq 1$ the variables $t_{n}$ and $\bar{t}_{n}$ are the holomorphic and antiholomorphic harmonic moments of $\mathbb{C}\setminus D$. The hierarchy possesses a dispersionless $\tau$-function whose logarithm $\mathcal{F}=\log\tau$ plays the role of a free energy, and in a standard normalization one has
$H_{mn}=r^{m+n}\frac{\partial^{2}\mathcal{F}}{\partial t_{m}\,\partial\bar{t}_{n}}.$
Thus $(H_{mn})$ is the scale-invariant mixed Hessian of the dispersionless free energy.
The same coefficients may be interpreted in several equivalent languages. In Laplacian growth and Hele–Shaw dynamics, $(H_{mn})$ records the second-order coupling between moment deformations. In the dispersionless Toda hierarchy, it is the matrix of mixed second derivatives of $\log\tau$. In Bergman-kernel language, one has, with our normalization,
(1.2) $\pi B_{\mathbb{C}\setminus D}(z,z^{\prime})=-\partial_{z}\partial_{\bar{z}^{\prime}}\log\!\left(1-\frac{1}{w(z)\,\overline{w(z^{\prime})}}\right),$
so the same data may be recovered from the exterior Bergman kernel in conformal coordinates [3, 11]. In the normal matrix model, $\mathcal{F}$ is the planar free energy, or equivalently the logarithmic energy of the equilibrium measure on $D$ [25, 23], and $(H_{mn})$ may be viewed as a mixed susceptibility matrix with respect to moment deformations.
The main contribution of the paper is to determine the spectral structure of $(H_{mn})$ near criticality in a family for which the inverse map can be controlled explicitly. The question is which part of the spectrum becomes singular near criticality, and whether that transition is governed by analytic criticality of the inverse branch or by the later geometric loss of univalence. This question is genuinely different from the classical holomorphic Grunsky theory, whose operator bounds are tied directly to univalence and quasiconformal extension [19, 5]. In the mixed sector studied here, the first instability is triggered already at the analytic threshold and concentrates into one logarithmically diverging stiff mode per symmetry sector, that is, an eigen-direction whose eigenvalue diverges logarithmically, while the remaining sectorial spectrum stays bounded. The resulting separation between analytic and geometric criticality is invisible in a purely holomorphic framework.
Relation to prior work. The present problem should also be distinguished from spectral and determinantal questions for the classical Grunsky operator and matrix, including Fredholm-eigenvalue problems for quasicircles and recent asymptotics for truncated Grunsky operators [21, 22, 12]. Those works concern the holomorphic Grunsky data, or Fredholm determinants built from them. Here the basic object is instead the mixed Hessian $(H_{mn})$, equivalently the mixed logarithmic kernel (1.1); the relevant transition is therefore not the saturation of a bounded holomorphic operator norm or a truncated determinant asymptotic, but a logarithmic rank-one spike in the mixed sector at the earlier analytic threshold $\zeta_{c}$.
With this motivation in hand, we now restrict to a family in which the analysis can be carried out explicitly. We study the $s$-fold symmetric one-harmonic family
(1.3) $f(w)=rw+aw^{1-s},\qquad s\geq 2,\quad r>0,\quad a\in\mathbb{R},$
which is the simplest family in which the analytic singularity of the inverse map and the geometric loss of univalence separate. Maps of the form (1.3) are standard in Laplacian growth and related free-boundary problems [9]; they already exhibit nontrivial geometric transitions, including loss of univalence and, for $s\geq 3$, boundary cusps at the geometric threshold. Because they depend on a single dimensionless shape parameter $\zeta:=a/r>0$, they provide a minimal and fully explicit model in which the spectral behavior of the mixed Hessian can be analyzed in detail. All rigorous results proved below concern the family (1.3); the corresponding extension to general polynomial conformal maps, under an isolated dominant square-root orbit hypothesis, is developed in the companion paper [2].
Writing $x:=r/z$ and $w(z)^{-1}=x\,U(x;\zeta)$, one finds that the inverse branch is governed by the algebraic equation
(1.4) $U=1+\zeta x^{s}U^{s}.$
This elementary relation already contains the two thresholds that control the problem. The first is the analytic threshold
$\zeta_{c}=\frac{(s-1)^{s-1}}{s^{s}},$
at which the dominant square-root singularity of the inverse branch reaches the unit circle in the $x$-plane. The second is the geometric threshold
$\zeta_{\mathrm{univ}}=\frac{1}{s-1},$
at which the conformal map ceases to be univalent and the boundary develops its first geometric singularity: for $s\geq 3$, semicubical cusps, and for $s=2$, the degenerate Joukowski segment. The problem is to determine which of these thresholds governs the first spectral instability of the mixed Hessian.
Our answer is that the relevant threshold is $\zeta_{c}$. The $s$-fold symmetry of (1.3) implies $H_{mn}=0$ unless $m\equiv n\pmod{s}$, so the Hessian decomposes into $s$ independent blocks. These blocks admit an exact positive Gram factorization in terms of Raney coefficients generated by (1.4). After passing to a suitable weighted realization, each symmetry block develops one logarithmically diverging eigenvalue, while the remaining sectorial spectrum stays bounded. Equivalently, there is one stiff mode per symmetry sector, so at most $s$ logarithmically diverging eigenvalues globally. The first spectral transition is therefore an analytic phenomenon of the inverse branch rather than a manifestation of geometric non-univalence.
A second theme of the paper concerns the intermediate regime
$\zeta_{c}<\zeta<\zeta_{\mathrm{univ}},$
where the conformal map remains univalent although the weighted compact operator realization used in the subcritical regime is no longer available. In this region the block-operator picture breaks down, but its scalar Gram building blocks remain well defined and continue analytically. More precisely, the block Gram weights in the subcritical operator theory are obtained from generating functions built from the squared Raney coefficients, and these scalar functions still admit analytic continuation beyond $\zeta_{c}$. We identify these continued functions as generalized hypergeometric objects, derive their resonant branch-point expansion, prove a Cauchy–Stieltjes representation, and, in the positive range $1\leq p\leq s$, obtain a Jacobi-matrix realization. Thus the continuation of these scalar Gram data captures the same analytic branch-point mechanism that produces the subcritical logarithmic instability, even after the weighted compact realization has ceased to exist. In particular, the singularity at $\zeta_{c}$ is not the onset of geometric breakdown: the two lateral boundary values of the continued scalar data remain finite at the univalence threshold.
1.1. Main results
The results of the paper fall into three groups. Theorem A gives the operator-theoretic description of the analytic threshold $\zeta_{c}$: for a fixed weight parameter $\beta>0$, each symmetry block splits into one logarithmically stiff direction and a bounded soft remainder. Theorem B identifies the scalar Gram data that survive beyond the weighted compact regime and continues them across the cut as generalized hypergeometric/Stieltjes objects. Finally, Proposition C shows that this analytic transition occurs strictly before the geometric loss of univalence at $\zeta_{\mathrm{univ}}$.
Fix a symmetry sector $q\in\{1,\dots,s\}$ and a weight parameter $\beta>0$. Let $H^{(q)}$ denote the intrinsic coefficient matrix of the $q$-th Hessian block, and let $\widetilde{G}^{(q)}(\zeta)$ be the corresponding weighted block Gram operator. The spectral analysis in this paper is carried out for these weighted realizations on the fixed Hilbert space $\ell^{2}(\mathbb{N}_{0})$, where $\mathbb{N}_{0}:=\{0,1,2,\dots\}$. The weighting is only an auxiliary Hilbert-space device: it does not alter the coefficient data and does not remove the critical instability. The weighted Gram and weighted Hessian blocks have the same nonzero spectrum, so the spectral conclusions may be stated equivalently for $\widetilde{G}^{(q)}(\zeta)$ or $\widetilde{H}^{(q)}(\zeta)$.
Theorem (A: Rank-one logarithmic instability).
Fix $q\in\{1,\dots,s\}$ and $\beta>0$. The weighted block Gram operator admits, as $\zeta\uparrow\zeta_{c}$, a rank-one logarithmic decomposition
$\widetilde{G}^{(q)}(\zeta)=L(\zeta)\,\widetilde{\bm{d}}^{(q)}\otimes\widetilde{\bm{d}}^{(q)*}+\widetilde{C}^{(q)}(\zeta),\qquad L(\zeta):=\log\Bigl(\frac{1}{1-\zeta^{2}/\zeta_{c}^{2}}\Bigr)\to+\infty,$
where $\widetilde{\bm{d}}^{(q)}\neq 0$ is independent of $\zeta$, while $\widetilde{C}^{(q)}(\zeta)$ remains uniformly bounded and converges in operator norm to a compact limit.
In particular, exactly one eigenvalue in the $q$-th sector diverges logarithmically,
$\mu_{1}^{(q)}(\zeta)=\|\widetilde{\bm{d}}^{(q)}\|_{\ell^{2}}^{2}\,L(\zeta)+O(1).$
All remaining sectorial eigenvalues stay bounded; more precisely, after removing the spike direction, each fixed soft eigenvalue converges to the corresponding eigenvalue of the compressed compact limit described in Proposition 4.8.
Theorem A is proved by combining the square-root singularity analysis of the inverse branch with the blockwise Gram representation of the Hessian. The logarithmic term is first isolated at the level of matrix entries, and the weighted realization for fixed $\beta>0$ then yields a compact operator framework in which the rank-one singular part separates from the bounded soft sector.
Remark 1.1 (Contrast with the Grunsky operator).
The classical Grunsky operator belongs to the holomorphic sector and is governed by the sharp bound $\|B\|\leq 1$, which is equivalent to univalence of the underlying map [19]. Accordingly, in the present one-parameter family, the holomorphic-sector transition is governed by the geometric threshold $\zeta_{\mathrm{univ}}$. By contrast, the mixed Hessian studied here transitions earlier, at the analytic threshold $\zeta_{c}$, and the nature of the transition is different: instead of the saturation of a bounded operator norm, one eigenvalue diverges logarithmically while the remaining spectrum stays bounded. This distinction between the holomorphic and mixed sectors is genuinely operator-theoretic and is not visible at the level of individual matrix entries.
The second principal result concerns the scalar quantities that underlie the block Gram construction itself and therefore remain meaningful beyond the weighted operator regime. For each $p\geq 1$, consider the generating function
(1.5) $\mathcal{G}_{p}(u):=\sum_{m\geq 0}R_{s,p}(m)^{2}\,u^{m},\qquad u=\zeta^{2}.$
Below $\zeta_{c}$, the scalar Gram weights entering the block operators are obtained from these functions by explicit Euler-type differential operators. The importance of $\mathcal{G}_{p}$ is therefore that they encode, in scalar form, the same branch-point mechanism that drives the logarithmic spectral instability. These functions continue analytically beyond the radius of convergence $u=\zeta_{c}^{2}$, even though the weighted compact operator picture no longer persists there.
Theorem (B: Analytic continuation of the scalar Gram data).
For each integer $p\geq 1$, the function $\mathcal{G}_{p}$ extends from its disk of convergence to a single-valued holomorphic function on $\mathbb{C}\setminus[\zeta_{c}^{2},\infty)$. It is a generalized hypergeometric function, and near the branch point $u=\zeta_{c}^{2}$ it admits a resonant expansion of the form
(1.6) $\mathcal{G}_{p}(u)=A(u)+B(u)\Bigl(1-\frac{u}{\zeta_{c}^{2}}\Bigr)^{2}\log\Bigl(1-\frac{u}{\zeta_{c}^{2}}\Bigr),$
with $A$ and $B$ analytic near $\zeta_{c}^{2}$. The scalar Gram weights are recovered from $\mathcal{G}_{p}$ by explicit Euler-type differential operators; these remove the quadratic prefactor and produce the logarithmic divergence at $u=\zeta_{c}^{2}$.
Theorem B is obtained from the explicit coefficient ratio of $R_{s,p}(m)^{2}$, which identifies $\mathcal{G}_{p}$ with a generalized hypergeometric function. The local expansion (1.6) is then derived by Frobenius analysis at the regular singular point $u=\zeta_{c}^{2}$. Moreover, the boundary values across the cut define a discontinuity density, and $\mathcal{G}_{p}$ admits a Cauchy–Stieltjes representation. In the positive range $1\leq p\leq s$, the corresponding Stieltjes measure is positive, which yields a Jacobi-matrix realization of $\mathcal{G}_{p}$ as a Weyl $m$-function.
The third principal result clarifies the relation between analytic and geometric criticality.
Proposition (C: Separation of thresholds).
For every $s\geq 2$,
(1.7) $\zeta_{c}<\zeta_{\mathrm{univ}}.$
Hence the logarithmic spectral instability of Theorem Theorem occurs while the conformal map is still univalent and the boundary is still smooth. Moreover, the analytically continued scalar data remain finite at $\zeta=\zeta_{\mathrm{univ}}$ (Propositions 6.3 and 6.5).
Thus the singularity of the Hessian at $\zeta_{c}$ is not a consequence of cusp formation. It is already encoded in the branch-point structure of the inverse map and therefore precedes geometric breakdown.
1.2. Organization of the paper
Section 2 analyzes the inverse conformal map, derives the Raney-number formulas, and identifies the square-root singularity at the analytic threshold. Section 3 introduces the scale-invariant Toda Hessian, proves the blockwise Gram factorization, and records the scalar Gram quantities that will later be continued across the threshold. Section 4 develops the weighted spectral theory in the subcritical regime $0<\zeta<\zeta_{c}$ and proves Theorem A. Section 5 studies the scalar continuation problem beyond $\zeta_{c}$, including the hypergeometric representation, the resonant expansion, the Cauchy–Stieltjes formula, and the Jacobi realization. Finally, Section 6 treats the geometric threshold $\zeta_{\mathrm{univ}}$, proves the strict separation (1.7), and shows that the continued scalar data remain finite at the univalence boundary. The appendices contain the uniform Raney asymptotics, the entrywise rank-one extraction, the hypergeometric continuation argument, and the computation of the branch-point coefficient.
2. Inversion of the conformal map
This section analyzes the inverse conformal map and extracts the information needed for the spectral analysis of the Hessian. It provides three inputs for later use: the functional equation for the inverse branch (Proposition 2.1), the explicit Taylor coefficients in terms of Raney numbers (Proposition 2.4), and the location and square-root nature of the dominant singularity (Lemma 2.2). The last point generates the universal $m^{-3/2}$ coefficient decay responsible for the logarithmic spectral divergence established in Section 4.
Throughout this manuscript we introduce the scale-free variable
(2.1) $x:=\frac{r}{z}.$
2.1. The generating function and its functional equation
We encode the inverse branch $w(z)\sim z/r$ as $z\to\infty$ by factoring out the leading linear growth:
(2.2) $w(z)=\frac{z}{r}\,U(x;\zeta)^{-1},$
where $U$ is a scalar function to be determined. The normalization $w(z)\sim z/r$ corresponds to $U(0;\zeta)=1$.
Proposition 2.1 (Functional equation for the inverse map).
Let $f(w)=rw+aw^{1-s}$ with $\zeta=a/r$, and let $w(z)$ be the inverse branch normalized by $w(z)\sim z/r$ at infinity. Define $U(x;\zeta)$ by (2.2). Then $U$ satisfies the algebraic functional equation
(2.3) $U(x;\zeta)=1+\zeta x^{s}U(x;\zeta)^{s},\qquad U(0;\zeta)=1.$
Proof.
Substituting (2.2) into the defining relation $z=rw+aw^{1-s}$ and using $x=r/z$ gives
$1=U^{-1}+\zeta x^{s}U^{s-1}.$
Multiplying by $U$ yields (2.3). ∎
Throughout we restrict to the one-parameter family with $\zeta:=a/r>0$. For the present one-parameter family, changing the sign of $a$ amounts to a rigid rotation of the image domain by $\pi/s$.
2.2. Square-root criticality
We now determine the local behavior of $U(x;\zeta)$ at its dominant singularity. This is the key analytic input for the spectral analysis: the square-root nature of the singularity produces the universal $m^{-3/2}$ coefficient asymptotics that drive the logarithmic divergence of the Hessian.
Lemma 2.2 (Critical point and local expansion).
The function $U(x;\zeta)$ defined by $U=1+\zeta x^{s}U^{s}$ is a power series in the combination $\zeta x^{s}$, with radius of convergence $\zeta_{c}=(s-1)^{s-1}/s^{s}$ in that variable. Equivalently, for fixed $\zeta>0$ its radius of convergence in $x$ is $(\zeta_{c}/\zeta)^{1/s}$. Its only singularity on $|\zeta x^{s}|=\zeta_{c}$ is at $\zeta x^{s}=\zeta_{c}$. At this point the function has a square-root branch point with the local expansion
(2.4) $U(x;\zeta)=\frac{s}{s-1}-\kappa\sqrt{1-\frac{\zeta x^{s}}{\zeta_{c}}}+O\!\left(1-\frac{\zeta x^{s}}{\zeta_{c}}\right),\qquad\zeta x^{s}\to\zeta_{c},$
where $\kappa=\sqrt{2s/(s-1)^{3}}>0$.
Proof.
Consider the implicit function $F(y,x;\zeta):=y-1-\zeta x^{s}y^{s}$. The Taylor branch $U(x;\zeta)$ is defined by $F(U,x;\zeta)=0$ with $U(0;\zeta)=1$. A singularity of this branch occurs where the implicit function theorem fails, i.e., where $\partial_{y}F=0$:
$F(U,x;\zeta)=0,\qquad\partial_{y}F(U,x;\zeta)=1-s\zeta x^{s}U^{s-1}=0.$
From the second equation, $\zeta x^{s}U^{s-1}=1/s$. Substituting into the first equation:
$U=1+\zeta x^{s}\,U^{s}=1+\frac{U}{s},$
which gives $U_{c}:=s/(s-1)$. Substituting back:
$\zeta_{c}=\frac{1}{s\,U_{c}^{s-1}}=\frac{1}{s}\left(\frac{s-1}{s}\right)^{s-1}=\frac{(s-1)^{s-1}}{s^{s}}.$
Since $U$ is algebraic, its singularities occur only at discriminant points of the polynomial $F(y,x;\zeta)=y-1-\zeta x^{s}y^{s}$ in the variable $y$. Equivalently, they occur exactly when the system
$F(y,x;\zeta)=0,\qquad\partial_{y}F(y,x;\zeta)=0$
has a common root. The computation above shows that the only nonzero critical value of the product $\zeta x^{s}$ is $\zeta_{c}$.
To determine the local behavior, we verify that the singularity is simple (i.e., $\partial_{yy}F\neq 0$ at the critical point):
$\partial_{yy}F=-s(s-1)\zeta x^{s}U^{s-2}=-\frac{(s-1)}{U_{c}}=-\frac{(s-1)^{2}}{s}\neq 0.$
Moreover, near any critical point with $\zeta x^{s}=\zeta_{c}$ and $y=U_{c}$, the Taylor expansion of $F$ has the form
$F(y,x;\zeta)=\frac{1}{2}\,\partial_{yy}F(U_{c},x;\zeta)\,(y-U_{c})^{2}-U_{c}^{s}\,(\zeta x^{s}-\zeta_{c})+O\!\bigl(|y-U_{c}|^{3}+|\zeta x^{s}-\zeta_{c}|\,|y-U_{c}|\bigr),$
because $\partial_{y}F(U_{c},x;\zeta)=0$ at criticality and the coefficient of $(\zeta x^{s}-\zeta_{c})$ is $U_{c}^{s}$. Since $\partial_{yy}F(U_{c},x;\zeta)\neq 0$, the Weierstrass preparation theorem gives a square-root branch point, hence (2.4). Matching the leading coefficients yields
$\frac{1}{2}\,\partial_{yy}F(U_{c},x;\zeta)\,\kappa^{2}+U_{c}^{s}\zeta_{c}=0.$
Using $\partial_{yy}F(U_{c},x;\zeta)=-(s-1)^{2}/s$ and $U_{c}^{s}\zeta_{c}=1/(s-1)$, we obtain
$\kappa^{2}=\frac{2s}{(s-1)^{3}},$
as claimed. ∎
Remark 2.3 (Geometry in the $z$-plane).
Returning to the original variable $z$ via $x=r/z$, the condition $|\zeta x^{s}|=\zeta_{c}$ becomes $|z|=r(\zeta/\zeta_{c})^{1/s}$. Thus the $s$ branch points of the inverse map $w(z)$ lie on a circle whose radius shrinks toward $r$ as $\zeta\uparrow\zeta_{c}$. When $\zeta=\zeta_{c}$, the branch points reach the circle $|z|=r$, which is the image of the unit circle under the leading term $z=rw$ of the conformal map. This is the geometric manifestation of the analytic threshold.
2.3. Raney numbers and coefficient formulas
The Taylor coefficients of powers of $U$ admit a classical closed form. We define these coefficients directly and then identify them with the Raney numbers.
For any $p\in\mathbb{Z}$, write
(2.5) $U(x;\zeta)^{p}=\sum_{n=0}^{\infty}R_{s,p}(n)\,(\zeta x^{s})^{n}.$
Proposition 2.4 (Raney numbers).
For all integers $s\geq 2$, $p\geq 1$, and $n\geq 0$,
(2.6) $R_{s,p}(n)=\frac{p}{sn+p}\binom{sn+p}{n}$.
For $p\geq 1$, these are the classical Raney numbers; the case $p=1$ gives the Fuss–Catalan numbers [20, 8].
Proof.
Set $U(t)=1+V(t)$ with $V(0)=0$. Then $V=t(1+V)^{s}$, which is the Lagrange form $V=t\Phi(V)$ with $\Phi(v)=(1+v)^{s}$. We apply the Lagrange–Bürmann inversion formula to the composite function $F(V)=(1+V)^{p}=U^{p}$. The standard coefficient extraction [10, 4] gives, for $n\geq 1$,
$t^{n}]F(V(t))=\frac{1}{n}[v^{n-1}]F^{\prime}(v)\Phi(v)^{n}.$
Since $F^{\prime}(v)=p(1+v)^{p-1}$ and $\Phi(v)^{n}=(1+v)^{sn}$, we obtain
$t^{n}]U(t)^{p}=\frac{p}{n}[v^{n-1}](1+v)^{sn+p-1}=\frac{p}{n}\binom{sn+p-1}{n-1}$.
The identity $\frac{p}{n}\binom{sn+p-1}{n-1}=\frac{p}{sn+p}\binom{sn+p}{n}$ yields (2.6) for $n\geq 1$. For $n=0$, we have $R_{s,p}(0)=U(0)^{p}=1$ (when $p\neq 0$), which agrees with (2.6). ∎
Remark 2.5 (Extension to $p\leq 0$).
The generating function $U(x;\zeta)^{p}$ is well-defined for all $p\in\mathbb{Z}$. For $p=0$, we have $U^{0}\equiv 1$, so $R_{s,0}(n)=\delta_{n,0}$. For $p<0$, the formula (2.6) can be extended using the generalized binomial coefficient, but requires separate treatment at values of $n$ where $sn+p=0$. Since all applications in this paper involve $p\geq 1$ (the index $p$ in the Hessian expansion), we restrict to positive $p$ in the main statement.
Corollary 2.6 (Convolution property).
For any integers $k\geq 1$ and $p_{1},\ldots,p_{k}\in\mathbb{Z}$, and any $m\geq 0$,
(2.7) $\sum_{\begin{subarray}{c}n_{1}+\cdots+n_{k}=m\\ n_{i}\geq 0\end{subarray}}\prod_{i=1}^{k}R_{s,p_{i}}(n_{i})=R_{s,p_{1}+\cdots+p_{k}}(m).$
Proof.
Multiply the generating series (2.5) and extract the coefficient of $(\zeta x^{s})^{m}$. ∎
We now record the expansions needed for the Hessian kernel in Section 3. The kernel involves the logarithm $\log(1-(w\bar{w}^{\prime})^{-1})$, which we will expand in powers of $w^{-1}$; hence we need the following.
Lemma 2.7 (Powers of the inverse map).
For each integer $p\geq 1$,
(2.8) $w(z)^{-p}=\left(\frac{r}{z}\right)^{p}\sum_{m=0}^{\infty}R_{s,p}(m)\,(\zeta x^{s})^{m}=\sum_{m=0}^{\infty}R_{s,p}(m)\,\zeta^{m}\,x^{p+ms}.$
In particular, the exponents of $x$ appearing in the expansion of $w(z)^{-p}$ are exactly $\{p+ms:m\geq 0\}$.
Proof.
From (2.2), $w^{-1}=(r/z)U=xU$, so $w^{-p}=x^{p}U^{p}$. Expanding $U^{p}$ via (2.5) gives (2.8). ∎
Remark 2.8 (Coefficient asymptotics).
The square-root singularity (2.4) implies, via standard transfer theorems [6], the universal large-$n$ asymptotics
(2.9) $R_{s,p}(n)=A_{s,p}\,n^{-3/2}\,\zeta_{c}^{-n}\bigl(1+O(n^{-1})\bigr),\qquad n\to\infty,$
with an explicit amplitude $A_{s,p}>0$ depending on $s$ and $p$. The exponent 3/2-3/2 is universal for simple algebraic singularities of square-root type. This is the only asymptotic information about the inverse map that enters the Hessian estimates in Section 4; the detailed derivation is given in Appendix A.
3. The Hessian of the $\tau$-function
This section constructs the main object of the paper: the mixed Hessian of $\log\tau$ with respect to the dispersionless Toda times. We work throughout in the scale-free variables introduced in Section 2, which removes the trivial dependence on the conformal radius and yields a Hessian that depends only on the shape parameter $\zeta=a/r$.
The section is organized as follows. We first recall the kernel representation due to Wiegmann–Zabrodin and use it to define the scale-invariant Hessian (Subsection 3.1). We then expand the kernel using the Raney machinery from Section 2, obtaining an exact Gram representation as a sum of rank-one operators (Subsection 3.3). Finally, we analyze the Gram weights and identify the analytic threshold $\zeta_{c}$ at which the naive unweighted Gram realization breaks down (Subsection 3.4).
3.1. Kernel representation and scale-invariant formulation
The dispersionless Toda hierarchy associates to a simply connected domain $D\subset\mathbb{C}$ a $\tau$-function whose logarithm $\mathcal{F}=\log\tau$ serves as a generating function for the harmonic moments. In the present paper, however, we use this framework only for the polynomial family (1.3). The general discussion from Section 1 is included only to place the problem in its geometric and integrable context.
For the maps (1.3) with $0<\zeta<\zeta_{\mathrm{univ}}$, the boundary is real-analytic and the Wiegmann–Zabrodin kernel identity applies. The mixed second derivatives of $\mathcal{F}$ with respect to the Toda times $(t_{m},\bar{t}_{n})$ are then expressed via the inverse conformal map $w(z)$ as follows [13]:
(3.1) $\frac{\partial^{2}\mathcal{F}}{\partial t_{m}\,\partial\bar{t}_{n}}=-mn\,[z^{-m}][\bar{z}^{\prime-n}]\log\!\left(1-\frac{1}{w(z)\,\overline{w(z^{\prime})}}\right),$
where $z$ and $z^{\prime}$ are treated as independent complex variables and the coefficient extraction $z^{-m}]$ refers to the Laurent expansion at infinity.
The right-hand side of (3.1) depends on the conformal radius $r$ through the normalization $w(z)\sim z/r$ at infinity. This dependence is inessential for spectral questions: it can be absorbed by passing to the scale-free variable $x=r/z$ introduced in Section 2. In terms of $x$, we have $w(z)^{-1}=x\,U(x;\zeta)$ by (2.2), and the kernel becomes
(3.2) ${K}(x,\bar{x}^{\prime}):=-\log\!\left(1-x\bar{x}^{\prime}\,U(x;\zeta)\,\overline{U(x^{\prime};\zeta)}\right).$
This expression depends on $(x,\bar{x}^{\prime},\zeta)$ alone, with no residual $r$-dependence.
Definition 3.1 (Scale-invariant Hessian).
The scale-invariant Hessian is the infinite matrix $H=(H_{mn})_{m,n\geq 1}$ defined by
(3.3) $H_{mn}:=mn\,[x^{m}][\bar{x}^{\prime n}]\,{K}(x,\bar{x}^{\prime}).$
The prefactor $mn$ arises from the coefficient extraction in (3.1) and ensures that $H$ has the correct homogeneity to act on mode sequences. Concretely, $H$ is the natural scale-invariant version of the Toda Hessian.
Remark 3.2 (Operator viewpoint and role of renormalization).
The scale-invariant Hessian $H=(H_{mn})_{m,n\geq 1}$ is naturally a positive semidefinite quadratic form on finitely supported sequences. Spectral statements require a compatible Hilbert realization, and the standard $\ell^{2}(\mathbb{N})$ topology is not stable at analytic criticality. The appropriate framework is the weighted Gram realization introduced later in Definition 4.1; all spectral statements in Sections 45 refer to these renormalized operators.
3.2. Expansion of the kernel and block structure
We now expand the kernel (3.2) using the Raney formalism from Section 2. The logarithm expands as
(3.4) ${K}(x,\bar{x}^{\prime})=\sum_{p=1}^{\infty}\frac{1}{p}\left(x\bar{x}^{\prime}\,U(x;\zeta)\,\overline{U(x^{\prime};\zeta)}\right)^{p}=\sum_{p=1}^{\infty}\frac{1}{p}\,x^{p}\bar{x}^{\prime p}\,U(x;\zeta)^{p}\,\overline{U(x^{\prime};\zeta)}{}^{p}$.
By Lemma 2.7, each power $U(x;\zeta)^{p}$ has the expansion
$U(x;\zeta)^{p}=\sum_{m=0}^{\infty}R_{s,p}(m)\,(\zeta x^{s})^{m}=\sum_{m=0}^{\infty}R_{s,p}(m)\,\zeta^{m}\,x^{sm},$
and similarly for the conjugate factor. Substituting into (3.4) and collecting powers of $x^{m}\bar{x}^{\prime n}$ yields:
Proposition 3.3 (Kernel coefficients).
The coefficient $x^{m}][\bar{x}^{\prime n}]{K}$ vanishes unless $m\equiv n\pmod{s}$. When this congruence holds,
(3.5) $x^{m}][\bar{x}^{\prime n}]{K}=\sum_{\begin{subarray}{c}p\geq 1\\ p\equiv m\pmod{s}\end{subarray}}\frac{1}{p}\,R_{s,p}(k)\,R_{s,p}(l)\,\zeta^{k+l}$
where k=(mp)/sk=(m-p)/s and l=(np)/sl=(n-p)/s.
Proof.
The monomial $x^{m}\bar{x}^{\prime n}$ arises from the $p$-th term in (3.4) when $m=p+sk$ and $n=p+sl$ for some $k,l\geq 0$. This requires $m\equiv p\equiv n\pmod{s}$. Summing over all contributing values of $p$ gives (3.5). ∎
The selection rule has an immediate structural consequence for the Hessian.
Corollary 3.4 (Block decomposition).
The scale-invariant Hessian decomposes as
(3.6) $H=\bigoplus_{q=1}^{s}H^{(q)},$
where $H^{(q)}$ is the restriction of $H$ to indices $m,n\equiv q\pmod{s}$. Explicitly, identifying the $q$-th block with an operator on $\ell^{2}(\mathbb{N}_{0})$ via the correspondence $m=q+js\leftrightarrow j$, the matrix elements are
(3.7) $H^{(q)}_{j_{1}j_{2}}=H_{q+j_{1}s,\,q+j_{2}s},\qquad j_{1},j_{2}\geq 0.$
All spectral statements in this paper are formulated and proved blockwise. The $s$-fold rotational symmetry of the conformal map thus reduces the spectral problem to $s$ independent components.
3.3. Gram representation
The factorized structure of (3.5), namely, a product of Raney coefficients, one depending on $m$ and one on $n$, suggests rewriting the Hessian as a sum of rank-one operators. This Gram representation is the key algebraic structure underlying the spectral analysis.
Proposition 3.5 (Gram representation).
Define vectors $\bm{v}^{(p)}=(v^{(p)}_{m})_{m\geq 1}$ by
(3.8) $v^{(p)}_{m}:=\begin{cases}\displaystyle\frac{m}{\sqrt{p}}\,R_{s,p}(k)\,\zeta^{k},&\text{if }m=p+ks\text{ for some }k\geq 0,\\[4.30554pt] 0,&\text{otherwise}.\end{cases}$
Then the scale-invariant Hessian admits the exact representation
(3.9) $H=\sum_{p=1}^{\infty}\bm{v}^{(p)}\otimes\bm{v}^{(p)*},$
where $\bm{v}\otimes\bm{v}^{*}$ denotes the rank-one operator $(\bm{v}\otimes\bm{v}^{*})_{mn}=v_{m}\overline{v_{n}}$.
Proof.
Combining Definition 3.1 with Proposition 3.3, we have for $m=p+ks$ and $n=p+ls$:
$H_{mn}=mn\sum_{p^{\prime}\equiv m}\frac{1}{p^{\prime}}R_{s,p^{\prime}}(k^{\prime})R_{s,p^{\prime}}(l^{\prime})\zeta^{k^{\prime}+l^{\prime}},$
where $k^{\prime}=(m-p^{\prime})/s$ and $l^{\prime}=(n-p^{\prime})/s$. On the other hand,
$\sum_{p=1}^{\infty}v^{(p)}_{m}\overline{v^{(p)}_{n}}=\sum_{p\equiv m\equiv n}\frac{mn}{p}R_{s,p}(k)R_{s,p}(l)\zeta^{k+l},$
with k=(mp)/sk=(m-p)/s and l=(np)/sl=(n-p)/s. These expressions coincide.
[…]
Remark 3.6 (Structure of the Gram representation).
The Gram form (3.9) makes positivity of $H$ manifest: it is a sum of positive semidefinite rank-one operators. More importantly, it separates two sources of complexity: (i) The individual contribution of each logarithmic mode $p$, encoded in the vector $\bm{v}^{(p)}$, and (ii) the collective mixing of modes within each symmetry sector, which determines the eigenvalue distribution.
The spectral behavior of $H$ is controlled by the interplay between these two effects. As we show next, the critical phenomenon at $\zeta=\zeta_{c}$ arises from the borderline summability of the Gram data, and not from any individual term becoming singular.
3.4. Gram weights and the analytic threshold
The Gram representation expresses $H$ as a superposition of rank-one contributions. Whether this sum defines a bounded operator on the naive unweighted $\ell^{2}$-space depends on the size of the individual terms, measured by their squared norms.
Definition 3.7 (Gram weights).
For $p\geq 1$, the $p$-th Gram weight is
(3.10) $\sigma_{p}(\zeta):=\|\bm{v}^{(p)}\|_{\ell^{2}}^{2}=\sum_{m=0}^{\infty}\frac{(p+ms)^{2}}{p}\,R_{s,p}(m)^{2}\,\zeta^{2m}.$
The Gram weights are the natural quantities controlling operator bounds for the Hessian. Their behavior as $\zeta$ varies reveals the analytic threshold.
Proposition 3.8 (Threshold for Gram weights).
For each fixed $p\geq 1$, one has $\sigma_{p}(\zeta)<\infty$ if and only if $\zeta<\zeta_{c}$. In particular, the analytic threshold for the scalar Gram weights is exactly $\zeta_{c}$.
Proof.
By Proposition A.1, the Raney coefficients satisfy $R_{s,p}(m)\sim A_{s,p}\,m^{-3/2}\,\zeta_{c}^{-m}$ as $m\to\infty$. Hence the summand in (3.10) has the borderline form
$\frac{(p+ms)^{2}}{p}R_{s,p}(m)^{2}\zeta^{2m}\sim\frac{s^{2}A_{s,p}^{2}}{p}\,m^{-1}\left(\frac{\zeta}{\zeta_{c}}\right)^{2m}.$
The factor $m^{-1}$ is summable exactly when $\zeta<\zeta_{c}$, and not summable at $\zeta=\zeta_{c}$. This proves the claim. ∎
Remark 3.9 (Nature of the threshold).
Proposition 3.8 shows that $\zeta_{c}$ is exactly the threshold at which the vectors $\bm{v}^{(p)}$ cease to belong to $\ell^{2}$. Hence the elementary Gram realization of this subsection is valid only for $\zeta<\zeta_{c}$. The coefficient matrix $H$, however, remains well defined throughout the geometric regime $\zeta<\zeta_{\mathrm{univ}}$, and the weighted framework of Section 4 is introduced to study the singular behavior near $\zeta_{c}$ on a fixed Hilbert scale.
The strict inequality $\zeta_{c}<\zeta_{\mathrm{univ}}$, proved later in Proposition 6.3, produces the intermediate regime $\zeta_{c}<\zeta<\zeta_{\mathrm{univ}}$ in which spectral criticality precedes geometric breakdown. The detailed analysis of these two regimes is carried out in Sections 4 and 5.
4. Subcritical phase ($0<\zeta<\zeta_{c}$)
This section develops the operator-theoretic picture in the analytic subcritical regime. Proposition 3.5 gives a canonical Gram factorization of each symmetry block of the Hessian. Although the unrenormalized Gram weights $\sigma_{p}(\zeta)$ diverge as $\zeta\uparrow\zeta_{c}$ (Proposition 3.8), one can renormalize the column modes by an explicit diagonal weight. The resulting compact positive operators exhibit a single logarithmically diverging eigenvalue, while the remaining soft spectrum stays bounded and admits a well-defined limiting description after compression.
4.1. Operator setting, block factorization, and weighted renormalization
Fix a symmetry sector $q\in\{1,\dots,s\}$, and write
$p_{j}:=q+js,\qquad j\in\mathbb{N}_{0},$
for the indices belonging to that block; see Corollary 3.4. Restricting the Gram vectors $\bm{v}^{(p)}$ from (3.8) to the $q$-th block, we obtain a synthesis map $V(\zeta)$, initially defined on finitely supported sequences, whose $j$-th column is the restriction of $\bm{v}^{(p_{j})}$. The block Hessian and block Gram operators are then
(4.1) $H^{(q)}(\zeta)=V(\zeta)V(\zeta)^{*},\qquad G^{(q)}(\zeta)=V(\zeta)^{*}V(\zeta).$
At this stage, $V(\zeta)$ is used only on finitely supported sequences, and (4.1) is an identity of matrix coefficients (equivalently, of quadratic forms on finitely supported vectors).
The difficulty is that the unweighted realization is not stable as $\zeta\uparrow\zeta_{c}$. Indeed, Proposition A.2 shows that the Raney coefficients satisfy
$R_{s,p}(m)\leq C_{s}\,p\,M^{p}\,\zeta_{c}^{-m}\,m^{-3/2},\qquad M:=\frac{s}{s-1}.$
Accordingly, the $j$-th column of $V(\zeta)$ carries the noncritical background growth $p_{j}M^{p_{j}}$ in the block index $j$, together with the borderline transfer tail $m^{-3/2}$ in the summation variable $m$. To obtain a compact operator on a fixed Hilbert space, one therefore rescales the columns by an explicit diagonal weight that removes this background growth but leaves the genuine logarithmic singularity untouched.
Definition 4.1 (Weighted realization).
Fix $\beta>0$, set $M:=\frac{s}{s-1}$, and define
(4.2) $w_{j}:=p_{j}^{\,\frac{3}{2}+\beta}M^{p_{j}},\qquad j\in\mathbb{N}_{0}.$
Let
(4.3) $\mathcal{H}_{\beta}:=\Bigl\{x=(x_{j})_{j\geq 0}:\,\|x\|_{\mathcal{H}_{\beta}}^{2}=\sum_{j\geq 0}|x_{j}|^{2}w_{j}^{2}<\infty\Bigr\},$
with inner product
$\langle x,y\rangle_{\mathcal{H}_{\beta}}:=\sum_{j\geq 0}x_{j}\overline{y_{j}}\,w_{j}^{2}.$
Let $\mathcal{W}:\mathcal{H}_{\beta}\to\ell^{2}(\mathbb{N}_{0})$ be the unitary map
$(\mathcal{W}x)_{j}=w_{j}x_{j}.$
The renormalized synthesis operator is
$\widetilde{V}(\zeta):=V(\zeta)\mathcal{W}^{-1}:\ell^{2}(\mathbb{N}_{0})\to\ell^{2}(\mathbb{N}_{0}),$
initially defined on finitely supported sequences and extended by continuity. Its associated weighted Gram and Hessian operators are
(4.4) $\widetilde{G}^{(q)}(\zeta):=\widetilde{V}(\zeta)^{*}\widetilde{V}(\zeta),\qquad\widetilde{H}^{(q)}(\zeta):=\widetilde{V}(\zeta)\widetilde{V}(\zeta)^{*}.$
The structure of the weights in (4.2) is dictated by the Raney asymptotics. The factor $M^{p_{j}}$ compensates the exponential dependence on the block index. The exponent $3/2$ matches the borderline tail $m^{-3/2}$, and the additional parameter $\beta>0$ provides a small margin that turns the residual $p_{j}^{-2}$ behavior into an absolutely summable sequence. In particular, the weighted realization is $\zeta$-independent.
Remark 4.2 (Intrinsic object versus weighted realization).
The coefficient matrix of the Hessian block is the intrinsic object. The weights (4.2) do not modify these coefficients and do not remove the critical instability; they only realize the same coefficient data in a fixed, $\zeta$-independent Hilbert scale in which the singular part becomes visible as a compact-plus-rank-one operator.
The singular direction is already present at the coefficient level and is encoded by the intrinsic amplitudes $d_{j}^{(q)}$; see (B.4). The vector
$\widetilde{\bm{d}}^{(q)}=\bigl(d_{j}^{(q)}/w_{j}\bigr)_{j\geq 0}\in\ell^{2}(\mathbb{N}_{0})$
is simply its realization in the weighted scale. Varying $\beta>0$ changes this realization, but not the underlying one-dimensional singular direction. Thus Theorems 4.6 and 4.7 are statements about the weighted realization for fixed $\beta>0$, whereas their intrinsic content is the existence of a single dominant singular direction in each symmetry sector.
The first consequence of this construction is compactness for every fixed subcritical parameter.
Proposition 4.3 (Subcritical compactness).
Fix $q\in\{1,\dots,s\}$ and $\beta>0$. For every $0<\zeta<\zeta_{c}$, the operator
$\widetilde{V}(\zeta):\ell^{2}(\mathbb{N}_{0})\to\ell^{2}(\mathbb{N}_{0})$
is Hilbert–Schmidt. Consequently, $\widetilde{G}^{(q)}(\zeta)$ and $\widetilde{H}^{(q)}(\zeta)$ are trace class, hence compact.
Proof.
By construction, the $j$-th column of $\widetilde{V}(\zeta)$ is
$\widetilde{V}(\zeta)e_{j}=\frac{\bm{v}^{(p_{j})}}{w_{j}},$
restricted to the $q$-block. Therefore
(4.5) $\|\widetilde{V}(\zeta)\|_{\mathrm{HS}}^{2}=\sum_{j\geq 0}\|\widetilde{V}(\zeta)e_{j}\|_{\ell^{2}}^{2}=\sum_{j\geq 0}\frac{\|\bm{v}^{(p_{j})}\|_{\ell^{2}}^{2}}{w_{j}^{2}}=\sum_{j\geq 0}\frac{\sigma_{p_{j}}(\zeta)}{w_{j}^{2}}.$
It is thus enough to bound $\sigma_{p}(\zeta)$ uniformly in $p$.
Using the closed form for $R_{s,p}(m)$ from Proposition 2.4, the Gram weights may be written as
(4.6) $\sigma_{p}(\zeta)=p\sum_{m\geq 0}\binom{sm+p}{m}^{2}\zeta^{2m},\qquad p\geq 1.$
We claim that there exists $C_{s}>0$ such that, for all $p\geq 1$ and $m\geq 1$,
(4.7) $\binom{sm+p}{m}\leq C_{s}\,M^{p}\,\zeta_{c}^{-m}\,m^{-1/2}.$
Indeed, set $N:=sm+p$, and K:=Nm=(s1)m+pK:=N-m=(s-1)m+p. A two-sided Stirling bound gives
$\binom{N}{m}\leq C\,\frac{N^{N+\frac{1}{2}}}{m^{m+\frac{1}{2}}K^{K+\frac{1}{2}}}=C\,\frac{1}{\sqrt{m}}\sqrt{\frac{N}{K}}\exp\bigl(N\log N-m\log m-K\log K\bigr),$
with an absolute constant $C>0$. Writing $t:=p/m\geq 0$, so that $N=m(s+t)$ and K=m(s1+t)K=m(s-1+t), the exponent becomes $m((s+t)\log(s+t)-(s-1+t)\log(s-1+t))$. By the convexity estimate used in Proposition A.2,
$(s+t)\log(s+t)-(s-1+t)\log(s-1+t)\leq\log(\zeta_{c}^{-1})+t\log M,$
hence
$\exp\bigl(N\log N-m\log m-K\log K\bigr)\leq\zeta_{c}^{-m}M^{p}.$
Moreover,
$\sqrt{\frac{N}{K}}=\sqrt{\frac{s+t}{s-1+t}}\leq\sqrt{\frac{s}{s-1}},$
uniformly in $t\geq 0$. This proves (4.7).
Now fix $0<\zeta<\zeta_{c}$ and set $\eta:=\zeta/\zeta_{c}\in(0,1)$. Combining (4.6) with (4.7), and separating the term $m=0$, we obtain
$\sigma_{p}(\zeta)=p+p\sum_{m\geq 1}\binom{sm+p}{m}^{2}\zeta^{2m}\leq p+p\,C_{s}^{2}M^{2p}\sum_{m\geq 1}m^{-1}\eta^{2m}.$
Since $\sum_{m\geq 1}m^{-1}\eta^{2m}=-\log(1-\eta^{2})<\infty$, this yields
(4.8) $\sigma_{p}(\zeta)\leq C_{s}^{\prime}(\zeta)\,p\,M^{2p},$
for some finite constant $C_{s}^{\prime}(\zeta)$ depending on $\zeta$, but not on $p$.
Finally, using $w_{j}^{2}=p_{j}^{3+2\beta}M^{2p_{j}}$ and (4.5),
$\frac{\sigma_{p_{j}}(\zeta)}{w_{j}^{2}}\leq C_{s}^{\prime}(\zeta)\,\frac{p_{j}M^{2p_{j}}}{p_{j}^{3+2\beta}M^{2p_{j}}}=C_{s}^{\prime}(\zeta)\,p_{j}^{-2-2\beta}.$
Because $p_{j}=q+js\sim sj$ and $\beta>0$, the series $\sum_{j\geq 0}p_{j}^{-2-2\beta}$ converges. Hence $\widetilde{V}(\zeta)$ is Hilbert–Schmidt. Therefore
$\widetilde{G}^{(q)}(\zeta)=\widetilde{V}(\zeta)^{*}\widetilde{V}(\zeta),\qquad\widetilde{H}^{(q)}(\zeta)=\widetilde{V}(\zeta)\widetilde{V}(\zeta)^{*},$
are trace class, and in particular compact. ∎
Lemma 4.4 (Isospectrality of the weighted block realizations).
For each $0<\zeta<\zeta_{c}$, the compact positive operators $\widetilde{H}^{(q)}(\zeta)$ and $\widetilde{G}^{(q)}(\zeta)$ have the same nonzero eigenvalues, counted with multiplicities.
Proof.
By Proposition 4.3, the operator $\widetilde{V}(\zeta):\ell^{2}(\mathbb{N}_{0})\to\ell^{2}(\mathbb{N}_{0})$ is bounded. Therefore the standard relation between $\widetilde{V}\widetilde{V}^{*}$ and $\widetilde{V}^{*}\widetilde{V}$ applies. If $\widetilde{G}^{(q)}(\zeta)x=\mu\,x$ with $\mu>0$, then $\widetilde{V}(\zeta)x\neq 0$ and
$\widetilde{H}^{(q)}(\zeta)\bigl(\widetilde{V}(\zeta)x\bigr)=\widetilde{V}(\zeta)\widetilde{V}(\zeta)^{*}\widetilde{V}(\zeta)x=\widetilde{V}(\zeta)\widetilde{G}^{(q)}(\zeta)x=\mu\,\widetilde{V}(\zeta)x.$
The converse implication is identical, with $\widetilde{V}(\zeta)^{*}$ in place of $\widetilde{V}(\zeta)$. Multiplicities are preserved by the same correspondence. ∎
Corollary 4.5 (Transfer from the weighted Gram block to the weighted Hessian block).
Fix $q\in\{1,\dots,s\}$, $\beta>0$, and $0<\zeta<\zeta_{c}$. Then $\widetilde{G}^{(q)}(\zeta)$ and $\widetilde{H}^{(q)}(\zeta)$ have the same nonzero eigenvalues, counted with multiplicities. In particular, every nonzero-eigenvalue statement proved below for $\widetilde{G}^{(q)}(\zeta)$ holds verbatim for $\widetilde{H}^{(q)}(\zeta)$.
Proof.
The first statement is exactly Lemma $4.$ The remaining assertion follows because all spectral conclusions in Sections 4.24.3 are formulated only in terms of the nonzero eigenvalues and their multiplicities. ∎
4.2. Rank-one logarithmic spike at $\zeta_{c}$
We now isolate the singular part of the weighted block Gram operator as $\zeta\uparrow\zeta_{c}$. The point is that, after the weighted realization of Section 4.1, the borderline divergence is carried by a single vector $\widetilde{\bm{d}}^{(q)}$, whereas the remaining part stays compact and converges in norm. This is the operator-theoretic core of Theorem A.
Theorem 4.6 (Rank-one decomposition).
Fix $q\in\{1,\dots,s\}$ and $\beta>0$. There exist a nonzero vector $\widetilde{\bm{d}}^{(q)}\in\ell^{2}(\mathbb{N}_{0})$, independent of $\zeta$, and a family of compact operators $\widetilde{C}^{(q)}(\zeta)$ on $\ell^{2}(\mathbb{N}_{0})$ such that, for every $0<\zeta<\zeta_{c}$,
(4.9) $\widetilde{G}^{(q)}(\zeta)=L(\zeta)\,\widetilde{\bm{d}}^{(q)}\otimes\widetilde{\bm{d}}^{(q)*}+\widetilde{C}^{(q)}(\zeta),\qquad L(\zeta):=\log\frac{1}{1-\zeta^{2}/\zeta_{c}^{2}}.$
Moreover,
$\sup_{0<\zeta<\zeta_{c}}\|\widetilde{C}^{(q)}(\zeta)\|<\infty,\qquad\widetilde{C}^{(q)}(\zeta)\to\widetilde{C}_{*}^{(q)}\quad\text{in operator norm as }\zeta\uparrow\zeta_{c},$
for some compact limit operator $\widetilde{C}_{*}^{(q)}$.
The vector $\widetilde{\bm{d}}^{(q)}$ is described explicitly in Appendix B. Its intrinsic, weight-independent amplitudes $d^{(q)}$ are defined in (B.4), and
$\widetilde{d}^{(q)}_{j}=\frac{d^{(q)}_{j}}{w_{j}}$
is their realization in the weighted scale.
Proof.
The detailed coefficient analysis is contained in Appendix B; here we only record the reduction that produces (4.9).
Entrywise extraction of the logarithmic term. By Lemma B.1, each matrix element of $\widetilde{G}^{(q)}(\zeta)$ is given by a single series in the summation index $m$. Lemma B.3 applies the uniform Raney expansion of Lemma A.3 to the tail of this series and controls the initial range by the global bound of Proposition A.2. As a result, if
$\eta:=\frac{\zeta}{\zeta_{c}}\in(0,1),$
and
$(K_{\eta})_{j_{1}j_{2}}:=\eta^{|j_{1}-j_{2}|}\,\widetilde{d}^{(q)}_{j_{1}}\widetilde{d}^{(q)}_{j_{2}},\qquad j_{1},j_{2}\in\mathbb{N}_{0},$
then
(4.10) $\widetilde{G}^{(q)}(\zeta)=L(\zeta)\,K_{\eta}+R^{(q)}(\zeta),$
where $R^{(q)}(\zeta)$ is Hilbert–Schmidt,
$\sup_{0<\zeta<\zeta_{c}}\|R^{(q)}(\zeta)\|_{\mathrm{HS}}<\infty,$
and
$R^{(q)}(\zeta)\to\widetilde{C}_{*}^{(q)}\qquad\text{in Hilbert--Schmidt norm as }\zeta\uparrow\zeta_{c}.$
In particular, $R^{(q)}(\zeta)$ is compact for each $\zeta$, and the convergence also holds in operator norm. We denote this operator-norm limit by $\widetilde{C}_{*}^{(q)}$; thus the $R_{*}^{(q)}$ of Appendix B and the $\widetilde{C}_{*}^{(q)}$ used in the main text are the same operator.
Removal of the residual Toeplitz factor. Set
$K_{1}:=\widetilde{\bm{d}}^{(q)}\otimes\widetilde{\bm{d}}^{(q)*}.$
The kernel $K_{\eta}$ differs from $K_{1}$ only by the Toeplitz damping factor $\eta^{|j_{1}-j_{2}|}$. By Lemma B.4,
(4.11) $L(\zeta)\,\|K_{\eta}-K_{1}\|\longrightarrow 0,\qquad\zeta\uparrow\zeta_{c}.$
Therefore
$L(\zeta)\,K_{\eta}=L(\zeta)\,K_{1}+L(\zeta)\,(K_{\eta}-K_{1}),$
and the second term is negligible at the operator level.
Final decomposition. Combining (4.10) with (4.11), we obtain
$\widetilde{G}^{(q)}(\zeta)=L(\zeta)\,K_{1}+\Bigl(R^{(q)}(\zeta)+L(\zeta)(K_{\eta}-K_{1})\Bigr).$
Define
$\widetilde{C}^{(q)}(\zeta):=R^{(q)}(\zeta)+L(\zeta)(K_{\eta}-K_{1}).$
Each $\widetilde{C}^{(q)}(\zeta)$ is compact, since $R^{(q)}(\zeta)$ is Hilbert–Schmidt and $K_{\eta}-K_{1}$ is Hilbert–Schmidt by Lemma B.4. Moreover,
$\sup_{0<\zeta<\zeta_{c}}\|\widetilde{C}^{(q)}(\zeta)\|\leq\sup_{0<\zeta<\zeta_{c}}\|R^{(q)}(\zeta)\|+\sup_{0<\zeta<\zeta_{c}}L(\zeta)\|K_{\eta}-K_{1}\|<\infty,$
and
$\widetilde{C}^{(q)}(\zeta)\to\widetilde{C}_{*}^{(q)}\qquad\text{in operator norm as }\zeta\uparrow\zeta_{c},$
because $R^{(q)}(\zeta)\to\widetilde{C}_{*}^{(q)}$ in Hilbert–Schmidt norm and $L(\zeta)(K_{\eta}-K_{1})\to 0$ in operator norm.
Since $K_{1}=\widetilde{\bm{d}}^{(q)}\otimes\widetilde{\bm{d}}^{(q)*}$, this is exactly (4.9). ∎
4.3. Spectral asymptotics
We now convert the rank-one decomposition of Theorem 4.6 into spectral information. The conclusion is that the singular part produces one stiff eigenvalue in each symmetry block, while the remaining spectrum stays bounded and converges to the compressed limit operator.
Theorem 4.7 (Spectral asymptotics at the threshold).
Fix $q\in\{1,\dots,s\}$ and $\beta>0$. Let
$\mu_{1}^{(q)}(\zeta)\geq\mu_{2}^{(q)}(\zeta)\geq\cdots\geq 0$
denote the eigenvalues of $\widetilde{G}^{(q)}(\zeta)$, counted with multiplicity, and let $\widetilde{\bm{d}}^{(q)}\neq 0$ be the spike vector from Theorem 4.6. Set
$\Gamma^{(q)}:=\|\widetilde{\bm{d}}^{(q)}\|_{\ell^{2}}^{2}.$
Then, as $\zeta\uparrow\zeta_{c}$,
(4.12) $\mu_{1}^{(q)}(\zeta)=\Gamma^{(q)}\,L(\zeta)+O(1),\qquad\sup_{0<\zeta<\zeta_{c}}\mu_{k}^{(q)}(\zeta)<\infty\quad(k\geq 2).$
In particular, in the $q$-th symmetry block exactly one eigenvalue diverges logarithmically, and this eigenvalue is simple for $\zeta$ sufficiently close to $\zeta_{c}$.
Moreover, if $\psi_{1}(\zeta)$ is a unit eigenvector associated with $\mu_{1}^{(q)}(\zeta)$, then after fixing its phase one has
(4.13) $\psi_{1}(\zeta)\longrightarrow\frac{\widetilde{\bm{d}}^{(q)}}{\|\widetilde{\bm{d}}^{(q)}\|_{\ell^{2}}}\qquad\text{in }\ell^{2}\quad\text{as }\zeta\uparrow\zeta_{c}.$
Finally,
(4.14) $\mu_{1}^{(q)}(\zeta)=\Gamma^{(q)}\,L(\zeta)+\Lambda_{\mathrm{fin}}^{(q)}+o(1),\qquad\zeta\uparrow\zeta_{c},$
where
$\Lambda_{\mathrm{fin}}^{(q)}:=\frac{\langle\widetilde{\bm{d}}^{(q)},\widetilde{C}_{*}^{(q)}\widetilde{\bm{d}}^{(q)}\rangle}{\|\widetilde{\bm{d}}^{(q)}\|_{\ell^{2}}^{2}}.$
Theorem 4.7 has a clear numerical signature: after weighted renormalization, exactly one eigenvalue in each symmetry block grows on the logarithmic scale $L(\zeta)$, while the remaining eigenvalues stay bounded. Figure 1 illustrates this separation for the block operator $\widetilde{G}^{(q)}(\zeta)$ in the sector $q=1$ for $s=3$ and $s=5$. The top row shows that the leading branch $\mu_{1}^{(q)}(\zeta)$ is asymptotically affine in $L(\zeta)$, whereas the bottom row confirms that after division by $L(\zeta)$ only the stiff branch survives and all soft branches tend to zero.
Figure 1. Logarithmic spectral asymptotics of the weighted Gram block $\widetilde{G}^{(q)}(\zeta)$ for $s=3,5$ in sector $q=1$ ($\beta=1$, $N=30$). Top row: the leading eigenvalue $\mu_{1}^{(q)}(\zeta)$ plotted against $L(\zeta)$, showing the asymptotically affine law $\mu_{1}^{(q)}(\zeta)=\Gamma^{(q)}\,L(\zeta)+O(1)$; the dashed line is a tail linear fit. Bottom row: the normalized eigenvalues $\mu_{k}^{(q)}(\zeta)/L(\zeta)$. The ratio $\mu_{1}^{(q)}(\zeta)/L(\zeta)$ approaches a positive constant, whereas $\mu_{k}^{(q)}(\zeta)/L(\zeta)\to 0$ for $k\geq 2$.
Proof.
Write $\widetilde{\bm{d}}:=\widetilde{\bm{d}}^{(q)},\ \widehat{\bm{d}}:=\widetilde{\bm{d}}/\|\widetilde{\bm{d}}\|_{\ell^{2}},\ P:=\widehat{\bm{d}}\otimes\widehat{\bm{d}}^{*}$. By Theorem 4.6,
(4.15) $\widetilde{G}^{(q)}(\zeta)=\Gamma^{(q)}\,L(\zeta)\,P+\widetilde{C}^{(q)}(\zeta),$
where each $\widetilde{C}^{(q)}(\zeta)$ is compact self-adjoint,
$K_{q}:=\sup_{0<\zeta<\zeta_{c}}\|\widetilde{C}^{(q)}(\zeta)\|<\infty,$
and $\widetilde{C}^{(q)}(\zeta)\to\widetilde{C}_{*}^{(q)}$ in operator norm as $\zeta\uparrow\zeta_{c}$.
Only one eigenvalue can diverge. If $x\perp\widetilde{\bm{d}}$ and $\|x\|=1$, then $Px=0$, hence
$\langle x,\widetilde{G}^{(q)}(\zeta)x\rangle=\langle x,\widetilde{C}^{(q)}(\zeta)x\rangle.$
Using the one-dimensional trial space $\operatorname{Span}\{\widetilde{\bm{d}}\}$ in the min–max characterization of $\mu_{2}^{(q)}(\zeta)$, we obtain
(4.16) $\mu_{2}^{(q)}(\zeta)\leq\sup_{\begin{subarray}{c}\|x\|=1\\ x\perp\widetilde{\bm{d}}\end{subarray}}\langle x,\widetilde{G}^{(q)}(\zeta)x\rangle\leq\|\widetilde{C}^{(q)}(\zeta)\|\leq K_{q}.$
Therefore every eigenvalue with index $k\geq 2$ remains uniformly bounded.
On the other hand, the Rayleigh quotient of $\widehat{\bm{d}}$ gives
$\mu_{1}^{(q)}(\zeta)\geq\langle\widehat{\bm{d}},\widetilde{G}^{(q)}(\zeta)\widehat{\bm{d}}\rangle=\Gamma^{(q)}\,L(\zeta)+\langle\widehat{\bm{d}},\widetilde{C}^{(q)}(\zeta)\widehat{\bm{d}}\rangle\geq\Gamma^{(q)}\,L(\zeta)-K_{q}.$
Since $L(\zeta)\to+\infty$, it follows that $\mu_{1}^{(q)}(\zeta)\to+\infty$ while $\mu_{k}^{(q)}(\zeta)$, $k\geq 2$, stay bounded. This proves (4.12), and the simplicity of $\mu_{1}^{(q)}(\zeta)$ for $\zeta$ close to $\zeta_{c}$ follows from the spectral gap
$\mu_{1}^{(q)}(\zeta)-\mu_{2}^{(q)}(\zeta)\to+\infty.$
The top eigenvector aligns with the spike direction. Let $\psi_{1}(\zeta)$ be a unit eigenvector corresponding to $\mu_{1}^{(q)}(\zeta)$. Applying IPI-P to the eigenvalue equation
$\widetilde{G}^{(q)}(\zeta)\psi_{1}(\zeta)=\mu_{1}^{(q)}(\zeta)\psi_{1}(\zeta)$
and using (IP)P=0(I-P)P=0, we obtain
$(I-P)\widetilde{C}^{(q)}(\zeta)\psi_{1}(\zeta)=\mu_{1}^{(q)}(\zeta)(I-P)\psi_{1}(\zeta).$
Hence
(4.17) $\|(I-P)\psi_{1}(\zeta)\|\leq\frac{\|\widetilde{C}^{(q)}(\zeta)\|}{\mu_{1}^{(q)}(\zeta)}\leq\frac{K_{q}}{\mu_{1}^{(q)}(\zeta)}=O\bigl(L(\zeta)^{-1}\bigr).$
If $a(\zeta):=\langle\psi_{1}(\zeta),\widehat{\bm{d}}\rangle$, then
$|a(\zeta)|^{2}=1-\|(I-P)\psi_{1}(\zeta)\|^{2}=1+O\bigl(L(\zeta)^{-2}\bigr).$
After fixing the phase so that $a(\zeta)\geq 0$, we obtain (4.13).
Asymptotics of the diverging eigenvalue. Since $\psi_{1}(\zeta)$ is normalized,
$\mu_{1}^{(q)}(\zeta)=\langle\psi_{1}(\zeta),\widetilde{G}^{(q)}(\zeta)\psi_{1}(\zeta)\rangle=\Gamma^{(q)}\,L(\zeta)\,|a(\zeta)|^{2}+\langle\psi_{1}(\zeta),\widetilde{C}^{(q)}(\zeta)\psi_{1}(\zeta)\rangle.$
By (4.17), $\Gamma^{(q)}\,L(\zeta)\,|a(\zeta)|^{2}=\Gamma^{(q)}\,L(\zeta)+o(1)$, while (4.13) and the operator-norm convergence of $\widetilde{C}^{(q)}(\zeta)$ imply
$\langle\psi_{1}(\zeta),\widetilde{C}^{(q)}(\zeta)\psi_{1}(\zeta)\rangle\to\langle\widehat{\bm{d}},\widetilde{C}_{*}^{(q)}\widehat{\bm{d}}\rangle=\frac{\langle\widetilde{\bm{d}},\widetilde{C}_{*}^{(q)}\widetilde{\bm{d}}\rangle}{\|\widetilde{\bm{d}}\|_{\ell^{2}}^{2}}.$
This is exactly (4.14). ∎
The remaining spectral data are governed by the compression of the limiting remainder to the orthogonal complement of the spike direction.
Proposition 4.8 (Convergence of the soft spectrum).
Fix $q\in\{1,\dots,s\}$ and $\beta>0$. Define
$\widehat{\bm{d}}^{(q)}:=\frac{\widetilde{\bm{d}}^{(q)}}{\|\widetilde{\bm{d}}^{(q)}\|_{\ell^{2}}},\qquad P:=\widehat{\bm{d}}^{(q)}\otimes\widehat{\bm{d}}^{(q)*},\qquad Q:=I-P,$
and let
$\widetilde{C}_{*,\perp}^{(q)}:=Q\,\widetilde{C}_{*}^{(q)}\,Q\big|_{(\widetilde{\bm{d}}^{(q)})^{\perp}}.$
Let
$\mu_{2,*}^{(q)}\geq\mu_{3,*}^{(q)}\geq\cdots\geq 0$
be the eigenvalues of $\widetilde{C}_{*,\perp}^{(q)}$, listed in nonincreasing order and counted with multiplicity. Then, for each fixed $n\geq 2$,
(4.18) $\mu_{n}^{(q)}(\zeta)\longrightarrow\mu_{n,*}^{(q)}\qquad\text{as }\zeta\uparrow\zeta_{c}.$
Proposition 4.8 identifies the limiting bounded spectrum that remains after the stiff direction is removed. This bounded part is naturally seen at the level of the compressed operator, not in the full diverging spectrum. Figure 2 illustrates this soft regime for $s=3$ and $s=5$: the upper panels show the convergence of the first soft branches as $\zeta\uparrow\zeta_{c}$, while the lower panels display near-critical finite-$\zeta$ soft spectral profiles across the symmetry sectors. Together with Figure 1, this makes the stiff/soft decomposition visually explicit.
Figure 2. Soft spectrum after removal of the stiff direction for $s=3,5$ ($\beta=1$, $N=40$). Top row: the soft branches $\mu_{2}^{(q)}(\zeta),\dots,\mu_{6}^{(q)}(\zeta)$ in sector $q=1$, plotted against $1/L(\zeta)$. Their flattening as $1/L(\zeta)\to 0$ shows that the soft branches remain bounded as $\zeta\uparrow\zeta_{c}$. Bottom row: finite-$\zeta$ snapshots of the soft branches for all sectors $q=1,\dots,s$ at $\zeta/\zeta_{c}=0.9999$.
Proof.
Write $\widetilde{\bm{d}}:=\widetilde{\bm{d}}^{(q)}$, $\widehat{\bm{d}}:=\widetilde{\bm{d}}/\|\widetilde{\bm{d}}\|_{\ell^{2}}$, $P:=\widehat{\bm{d}}\otimes\widehat{\bm{d}}^{*}$, and Q:=IPQ:=I-P. Let $\psi_{1}(\zeta)$ be a normalized eigenvector corresponding to the simple eigenvalue $\mu_{1}^{(q)}(\zeta)$, and define
$P_{\zeta}:=\psi_{1}(\zeta)\otimes\psi_{1}(\zeta)^{*},\qquad Q_{\zeta}:=I-P_{\zeta}.$
By (4.13), $\|P_{\zeta}-P\|\leq 2\|\psi_{1}(\zeta)-\widehat{\bm{d}}\|_{\ell^{2}}\to 0$, and therefore
(4.19) $\|P_{\zeta}-P\|\to 0,\qquad\|Q_{\zeta}-Q\|\to 0\qquad\text{as }\zeta\uparrow\zeta_{c}.$
Since $P_{\zeta}$ is the spectral projection of the simple top eigenvalue, the restriction of $\widetilde{G}^{(q)}(\zeta)$ to $\operatorname{Ran}Q_{\zeta}$ has eigenvalues precisely $\mu_{2}^{(q)}(\zeta)\geq\mu_{3}^{(q)}(\zeta)\geq\cdots$. It is therefore enough to prove norm convergence of the compressed operators:
(4.20) $\bigl\|Q_{\zeta}\widetilde{G}^{(q)}(\zeta)Q_{\zeta}-Q\widetilde{C}_{*}^{(q)}Q\bigr\|\longrightarrow 0.$
Using (4.15), we have
$Q_{\zeta}\widetilde{G}^{(q)}(\zeta)Q_{\zeta}=\Gamma^{(q)}\,L(\zeta)\,Q_{\zeta}PQ_{\zeta}+Q_{\zeta}\widetilde{C}^{(q)}(\zeta)Q_{\zeta}.$
Because $P=\widehat{\bm{d}}\otimes\widehat{\bm{d}}^{*}$ is rank one and $Q_{\zeta}=I-P_{\zeta}$, one has the exact identity
[…]
hence $\|Q_{\zeta}PQ_{\zeta}\|=\|Q_{\zeta}\widehat{\bm{d}}\|^{2}=1-|\langle\psi_{1}(\zeta),\widehat{\bm{d}}\rangle|^{2}=\|(I-P)\psi_{1}(\zeta)\|^{2}$. Therefore (4.17) implies
(4.21) $L(\zeta)\,\|Q_{\zeta}PQ_{\zeta}\|=L(\zeta)\,\|(I-P)\psi_{1}(\zeta)\|^{2}=O\bigl(L(\zeta)^{-1}\bigr)\longrightarrow 0.$
Next, $\|Q_{\zeta}\widetilde{C}^{(q)}(\zeta)Q_{\zeta}-Q\widetilde{C}^{(q)}(\zeta)Q\|\leq 2\|\widetilde{C}^{(q)}(\zeta)\|\,\|Q_{\zeta}-Q\|\to 0$ by (4.19), while
$\|Q\widetilde{C}^{(q)}(\zeta)Q-Q\widetilde{C}_{*}^{(q)}Q\|\leq\|\widetilde{C}^{(q)}(\zeta)-\widetilde{C}_{*}^{(q)}\|\to 0$
by Theorem 4.6. Together with (4.21), this yields (4.20).
Both $Q_{\zeta}\widetilde{G}^{(q)}(\zeta)Q_{\zeta}$ and $\widetilde{C}_{*,\perp}^{(q)}$ are compact self-adjoint operators. Hence (4.20) and the min–max principle imply convergence of every fixed ordered eigenvalue, which is exactly (4.18). ∎
Remark 4.9 (Soft modes).
Proposition 4.8 shows that, after the unique stiff direction $\widetilde{\bm{d}}^{(q)}$ is removed, the remaining spectrum has a finite limit as $\zeta\uparrow\zeta_{c}$. Thus the bounded eigenvalues
$\mu_{2}^{(q)}(\zeta),\mu_{3}^{(q)}(\zeta),\dots$
are governed at leading order by the compressed limit operator $\widetilde{C}_{*,\perp}^{(q)}$. In particular, the logarithmic divergence of Theorem 4.7 is entirely carried by the rank-one spike.
5. Continuation of scalar Gram data beyond the analytic threshold
5.1. Intrinsic Hessian, weighted realization, and continued scalar data
Three related objects must be kept distinct beyond the analytic threshold. First, the intrinsic Hessian coefficients $H_{mn}$ and the corresponding unweighted block coefficients are attached to the conformal map itself and therefore remain meaningful as long as the geometric map exists. Second, the weighted block operators $\widetilde{G}^{(q)}(\zeta)$ from Section 4, constructed for a fixed choice of $\beta>0$, are a fixed-space Hilbert realization of the subcritical theory. Their role is to separate the logarithmically singular rank-one part from the bounded soft remainder. Third, the scalar quantities $\mathcal{G}_{p}$, $\sigma_{p}$, and $\rho_{p}$ encode the same branch-point information at the coefficient level and are the objects that can be continued beyond the weighted operator regime.
The weighted compact-operator framework of Section 4 is inherently subcritical. As $\zeta\uparrow\zeta_{c}$, the scalar Gram weights $\sigma_{p}(\zeta)$ acquire the borderline logarithmic divergence isolated in Theorem 4.6. Accordingly, the fixed-space weighted block realization used in Section 4 is not continued past $\zeta_{c}$: no bounded continuation of the compact operators $\widetilde{G}^{(q)}(\zeta)$ on the same weighted Hilbert space is asserted here.
What does continue is the scalar data built from the squared Raney coefficients. The purpose of this section is to isolate those scalar quantities, continue them analytically beyond the radius of convergence, and show that they still encode the branch-point mechanism responsible for the subcritical spectral instability. Thus this is a scalar continuation theory, not an extension of the weighted operator picture beyond $\zeta_{c}$. Since the geometric map remains univalent up to the larger threshold $\zeta_{\mathrm{univ}}$; see Proposition 6.1, the interval
$\zeta_{c}<\zeta<\zeta_{\mathrm{univ}}$
should be understood as a regime in which the fixed-space weighted spectral realization breaks down before geometric univalence is lost.
The argument has three parts. First, we introduce the scalar functions $\mathcal{G}_{p}(u)$, where $u=\zeta^{2}$, and show that they are generalized hypergeometric functions $_{2s}F_{2s-1}$ with parametric excess $\gamma=2$; see Proposition 5.4 and Corollary 5.5. This yields a canonical analytic continuation to the slit plane
$\mathbb{C}\setminus[\zeta_{c}^{2},\infty).$
Second, we analyze the singular point $u=\zeta_{c}^{2}$ and prove the resonant local expansion
$\mathcal{G}_{p}(u)=A(u)+B(u)\Bigl(1-\frac{u}{\zeta_{c}^{2}}\Bigr)^{2}\log\!\Bigl(1-\frac{u}{\zeta_{c}^{2}}\Bigr),$
with $A$ and $B$ analytic and $B(\zeta_{c}^{2})<0$; see Theorem 5.9. The Euler operator that reconstructs the scalar Gram weights from $\mathcal{G}_{p}$ removes the quadratic prefactor and thereby converts this mild resonant singularity into the logarithmic divergence found in Section 4; see Corollary 5.10. Third, the branch-cut discontinuity of $\mathcal{G}_{p}$ gives a Cauchy–Stieltjes representation; see Proposition 5.16. In the range $1\leq p\leq s$, the representing measure is positive, and $\mathcal{G}_{p}$ becomes the Weyl $m$-function of a bounded Jacobi operator $J_{p}$; see Proposition 5.19 and Theorem 5.21.
Throughout this section we work with the continuation variable $u=\zeta^{2}$. Later, in the Stieltjes/Jacobi interpretation, we shall also use the reciprocal spectral variable $t=u^{-1}$. To avoid confusion, we reserve $\rho_{p}(u)$ for the jump density of the continued scalar Gram weight and $\varrho_{p}(t)$ for the density of the representing measure of $\mathcal{G}_{p}$. For a function $f$ defined on the slit plane, we write
$\operatorname{Disc}f(u):=f(u+i0)-f(u-i0),\qquad u\in(\zeta_{c}^{2},\infty),$
for the discontinuity across the branch cut.
5.2. From Gram weights to scalar generating functions
We begin by separating the genuinely analytic part of the Gram weights from the polynomial prefactor coming from the block index. Recall from Definition 3.7 that
(5.1) $\sigma_{p}(\zeta)=\|\bm{v}^{(p)}\|^{2}=\sum_{m=0}^{\infty}\frac{(p+ms)^{2}}{p}\,R_{s,p}(m)^{2}\,\zeta^{2m}.$
For $\zeta<\zeta_{c}$ this series converges, while for $\zeta\geq\zeta_{c}$ it diverges. The divergent behavior is entirely encoded in the squared Raney coefficients, and it is therefore natural to strip away the explicit polynomial factor $(p+ms)^{2}/p$.
Definition 5.1 (Squared Raney generating function).
For $p\geq 1$, define
(5.2) $\mathcal{G}_{p}(u):=\sum_{m=0}^{\infty}R_{s,p}(m)^{2}\,u^{m},\qquad u\in\mathbb{C},$
initially in the disk $|u|<\zeta_{c}^{2}$.
The point of Definition 5.1 is that the original Gram weight is recovered from $\mathcal{G}_{p}$ by a fixed second-order Euler operator. Thus the continuation problem for $\sigma_{p}$ reduces to the continuation problem for $\mathcal{G}_{p}$.
Lemma 5.2 (Euler operator identity).
For $|u|<\zeta_{c}^{2}$,
(5.3) $\sigma_{p}(\zeta)=\frac{1}{p}\Bigl(p+s\,u\frac{d}{du}\Bigr)^{\!2}\mathcal{G}_{p}(u)\Big|_{u=\zeta^{2}}.$
Proof.
For each monomial $u^{m}$ one has
$\Bigl(p+s\,u\frac{d}{du}\Bigr)u^{m}=(p+sm)u^{m},$
hence
$\frac{1}{p}\Bigl(p+s\,u\frac{d}{du}\Bigr)^{\!2}u^{m}=\frac{(p+sm)^{2}}{p}\,u^{m}.$
Applying this termwise to the absolutely convergent series (5.2) in the disk $|u|<\zeta_{c}^{2}$, and then setting $u=\zeta^{2}$, gives (5.3). ∎
Formula (5.3) is the basic bridge between the subcritical operator theory and the continuation theory developed below. The operator picture detects the singularity through the logarithmic growth of $\sigma_{p}(\zeta)$, while the scalar continuation problem is governed by the analytic structure of $\mathcal{G}_{p}(u)$ at the branch point $u=\zeta_{c}^{2}$.
5.3. Hypergeometric structure and analytic continuation
We next identify the analytic structure of the scalar generating function $\mathcal{G}_{p}$ from Definition 5.1. The first point is that its radius of convergence is exactly the analytic threshold $u=\zeta_{c}^{2}$, which is already encoded in the large-$m$ asymptotics of the Raney coefficients.
Proposition 5.3 (Radius of convergence).
The series (5.2) has radius of convergence $\zeta_{c}^{2}$. More precisely, as $m\to\infty$,
(5.4) $R_{s,p}(m)^{2}=\frac{C_{s,p}}{m^{3}}\,\zeta_{c}^{-2m}\bigl(1+o(1)\bigr),$
where $C_{s,p}>0$ depends only on $(s,p)$.
Proof.
By Proposition A.1, equivalently by the squared asymptotic formula (A.2) from Appendix A, one has
$R_{s,p}(m)^{2}\sim C_{s,p}\,m^{-3}\,\zeta_{c}^{-2m}.$
This is exactly (5.4), and the radius of convergence is therefore $\zeta_{c}^{2}$. ∎
The representation (5.2) is local in $u$. To continue $\mathcal{G}_{p}$ beyond $|u|<\zeta_{c}^{2}$, one needs a closed analytic model. This is provided by a generalized hypergeometric expression, whose existence ultimately reflects the fact that the coefficient ratio $R_{s,p}(m+1)^{2}/R_{s,p}(m)^{2}$ is rational in $m$.
Proposition 5.4 (Hypergeometric representation).
For $|u|<\zeta_{c}^{2}$, one has
(5.5) $\mathcal{G}_{p}(u)={}_{2s}F_{2s-1}\!\left(\begin{matrix}\alpha_{1},\ldots,\alpha_{2s}\\ \beta_{1},\ldots,\beta_{2s-1}\end{matrix}\Big|\,\zeta_{c}^{-2}u\right),$
where the parameter multisets are
(5.6) $\{\alpha_{i}\}=2\times\left\{\frac{p+k}{s}:k=0,\ldots,s-1\right\},\qquad\{\beta_{j}\}=\{1\}\ \cup\ 2\times\left\{\frac{p+l}{s-1}:l=1,\ldots,s-1\right\}.$
Proof.
See Appendix C. ∎
Corollary 5.5 (Parametric excess).
For the hypergeometric data (5.5)–(5.6),
$\gamma:=\sum_{j}\beta_{j}-\sum_{i}\alpha_{i}=2$
for every $s\geq 2$ and $p\geq 1$.
Proof.
A direct computation gives
$\sum_{i}\alpha_{i}=2\sum_{k=0}^{s-1}\frac{p+k}{s}=2p+(s-1),$
while
$\sum_{j}\beta_{j}=1+2\sum_{l=1}^{s-1}\frac{p+l}{s-1}=1+2p+s.$
Hence $\gamma=2$. ∎
The positivity of the parametric excess has two immediate consequences. First, the branch point of $\mathcal{G}_{p}$ occurs at $\xi:=\zeta_{c}^{-2}u=1$, that is, at $u=\zeta_{c}^{2}$. Second, the hypergeometric differential equation is of resonant type there. In particular, $\gamma=2$ forces a logarithmic term with prefactor $(1-\xi)^{2}$, rather than a pure algebraic singularity.
Lemma 5.6 (Reduced hypergeometric data after cancellation).
Let $c_{p}$ denote the number of cancelled common upper and lower parameters in (5.5), counted with multiplicity, and set
$q_{p}:=2s-1-c_{p}.$
After those cancellations, the germ of $\mathcal{G}_{p}$ at $u=0$ is represented by a reduced generalized hypergeometric function of type $_{q_{p}+1}F_{q_{p}}$ with the same parametric excess $2$. In particular, the reduced equation is nontrivial and has finite singular points only at $\xi=0$ and $\xi=1$, where $\xi=\zeta_{c}^{-2}u$.
Proof.
Cancelling a common upper and lower parameter removes the same quantity from $\sum_{i}\alpha_{i}$ and $\sum_{j}\beta_{j}$, so the parametric excess is unchanged. Since Corollary 5.5 gives excess $2$ before cancellation, the reduced equation also has excess $2$.
If the reduction were of type $_{1}F_{0}$, its unique upper parameter would have to equal 2-2 in order to have excess $2$, which is impossible because every surviving upper parameter comes from the positive list (5.6). Thus $q_{p}\geq 1$, so the reduced form is genuinely of type $_{q_{p}+1}F_{q_{p}}$. For such equations the only finite singular points are 0 and $1$ in the variable $\xi$; see [18, §16.8, §16.11]. ∎
Corollary 5.7 (Analytic continuation).
The function $\mathcal{G}_{p}$ extends to a single-valued holomorphic function on the slit plane
$\mathbb{C}\setminus[\zeta_{c}^{2},\infty).$
Proof.
By Lemma 5.6, after cancellation of common upper and lower parameters the germ at $u=0$ is represented by a reduced generalized hypergeometric function of type $_{q_{p}+1}F_{q_{p}}$ with the same parametric excess $2$. For that reduced hypergeometric differential equation, the only finite singular points in the variable $\xi=\zeta_{c}^{-2}u$ are $\xi=0$ and $\xi=1$; equivalently, the only finite singular points in the variable $u$ are $u=0$ and $u=\zeta_{c}^{2}$; see [18, §16.8, §16.11]. Hence the germ at the origin continues uniquely along every path in $\mathbb{C}\setminus[\zeta_{c}^{2},\infty)$, which is simply connected. This yields a single-valued holomorphic continuation of $\mathcal{G}_{p}$ to the slit plane. ∎
5.4. Resonant expansion at the branch point
We now analyze the singularity of $\mathcal{G}_{p}$ at the endpoint $u=\zeta_{c}^{2}$ of the disk of convergence. Because the parametric excess is the integer $\gamma=2$, the local Frobenius basis at the branch point is resonant, and the singular term is of the form $(1-u/\zeta_{c}^{2})^{2}\log(1-u/\zeta_{c}^{2})$.
Lemma 5.8 (Reduced local model at the branch point).
Let $\xi=\zeta_{c}^{-2}u$, and write the germ of $\mathcal{G}_{p}$ at $\xi=0$ in its reduced hypergeometric form $_{q_{p}+1}F_{q_{p}}$, obtained by cancelling any common upper and lower parameters in (5.5). Then $\xi=1$ is a regular singular point of the reduced differential equation. Its local exponents are
$\{0,2\}\quad\text{if }q_{p}=1,\qquad 0,1,\dots,q_{p}-1,2\quad\text{if }q_{p}\geq 2.$
In particular, the exponent 0 is simple, while the exponent $1$ is either absent or simple. Hence no terms of the form $\log(1-\xi)$ or $(1-\xi)\log(1-\xi)$ can occur in the local continuation of the germ.
Proof.
By Lemma 5.6, after cancellation $\mathcal{G}_{p}$ satisfies a reduced generalized hypergeometric equation of type $_{q_{p}+1}F_{q_{p}}$ with parametric excess $\gamma=2$. For such an equation, the local exponents at $\xi=1$ are $0,1,\dots,q_{p}-1,\gamma$; see [18, §16.8]. Substituting $\gamma=2$ gives the stated list. The exponent 0 is therefore simple, and the exponent $1$ is either absent or simple. Consequently, the local continuation of the germ selected by the Taylor series at $\xi=0$ cannot contain logarithmic terms of order $(1-\xi)^{0}$ or $(1-\xi)^{1}$. The first order at which a logarithmic sector may occur is thus $(1-\xi)^{2}$. ∎
Theorem 5.9 (Resonant expansion).
There exist functions $A(u)$ and $B(u)$, analytic in a neighborhood of $u=\zeta_{c}^{2}$, such that
(5.7) $\mathcal{G}_{p}(u)=A(u)+B(u)\Bigl(1-\frac{u}{\zeta_{c}^{2}}\Bigr)^{2}\log\Bigl(1-\frac{u}{\zeta_{c}^{2}}\Bigr),$
as $u\to\zeta_{c}^{2}$ in $\mathbb{C}\setminus[\zeta_{c}^{2},\infty)$. Moreover,
$B(\zeta_{c}^{2})<0$
for all $s\geq 2$ and $p\geq 1$.
Proof.
Set $\xi:=u/\zeta_{c}^{2}$. By Corollary 5.7, the germ defined by the Taylor series at $\xi=0$ extends holomorphically to the slit plane $\mathbb{C}\setminus[1,\infty)$. By Lemma 5.6, after cancellation of any common upper and lower parameters in (5.5), this germ is a reduced generalized hypergeometric function of type $_{q_{p}+1}F_{q_{p}}$ with parametric excess
$\gamma=2.$
By Lemma 5.8, the local exponents at $\xi=1$ are 0 and $1$ with no multiplicity, so logarithmic terms of order $(1-\xi)^{0}$ and $(1-\xi)^{1}$ are excluded. For a reduced generalized hypergeometric equation of type $_{q_{p}+1}F_{q_{p}}$ with positive integer parametric excess $m$, the connection formulas at $\xi=1$ on the slit plane imply that the continued germ has a local representation consisting of an analytic part plus one resonant term of the form $(1-\xi)^{m}\log(1-\xi)$ with analytic coefficient; see [18, §16.8(ii), §16.11]. Since here $m=\gamma=2$, it follows that
$\mathcal{G}_{p}(\zeta_{c}^{2}\xi)=A_{0}(\xi)+B_{0}(\xi)(1-\xi)^{2}\log(1-\xi),$
where $A_{0}$ and $B_{0}$ are analytic near $\xi=1$. The exponent analysis from Lemma 5.8 shows that no lower-order logarithmic terms occur.
Returning to the variable $u=\zeta_{c}^{2}\xi$, we obtain (5.7) with
$A(u):=A_{0}(u/\zeta_{c}^{2}),\qquad B(u):=B_{0}(u/\zeta_{c}^{2}).$
The coefficient $B(\zeta_{c}^{2})$ is computed explicitly in Appendix D. More precisely, Proposition D.1 yields
$B(\zeta_{c}^{2})<0$
for all $s\geq 2$ and $p\geq 1$. Hence the logarithmic sector is nontrivial and starts precisely at order $(1-\xi)^{2}\log(1-\xi)$. ∎
The singularity in (5.7) is mild: the function $\mathcal{G}_{p}(u)$ itself stays finite at $u=\zeta_{c}^{2}$. The logarithmic divergence appears only after one returns to the Gram weights via the Euler operator from Lemma 5.2. In this sense, the scalar continuation problem is strictly softer than the operator-theoretic one.
Corollary 5.10 (Mechanism of Gram-weight divergence).
As $\zeta\uparrow\zeta_{c}$, equivalently $u=\zeta^{2}\uparrow\zeta_{c}^{2}$, the Gram weights satisfy
$\sigma_{p}(\zeta)=\widetilde{A}_{p}+\frac{2s^{2}}{p}\,B(\zeta_{c}^{2})\,\log\Bigl(1-\frac{\zeta^{2}}{\zeta_{c}^{2}}\Bigr)+o(1),$
where $\widetilde{A}_{p}\in\mathbb{R}$, and $B(\zeta_{c}^{2})<0$ is the coefficient from Theorem 5.9. In particular,
$\sigma_{p}(\zeta)\to+\infty$
logarithmically, in agreement with Proposition 3.8.
Proof.
Apply Lemma 5.2 to the expansion (5.7), with $u=\zeta^{2}$, and write $w:=1-u/\zeta_{c}^{2}$. Since $A(u)$ and $B(u)$ are analytic at $u=\zeta_{c}^{2}$, one has
$\mathcal{G}_{p}(u)=A(u)+B(\zeta_{c}^{2})\,w^{2}\log w+O(w^{3}\log w).$
The Euler operator $\bigl(p+s\,u\frac{d}{du}\bigr)^{2}$ preserves analyticity, so the contribution of $A(u)$ is a constant plus $o(1)$, while $O(w^{3}\log w)$ contributes only $o(1)$. For the singular term,
(5.8) $\Bigl(p+s\,u\frac{d}{du}\Bigr)^{2}\!\bigl[w^{2}\log w\bigr]=2s^{2}\log w+3s^{2}+O\bigl(w\log w\bigr),\qquad w\downarrow 0.$
Hence
$\sigma_{p}(\zeta)=\widetilde{A}_{p}+\frac{2s^{2}}{p}\,B(\zeta_{c}^{2})\log w+o(1),$
for some $\widetilde{A}_{p}\in\mathbb{R}$. Since $w=1-\zeta^{2}/\zeta_{c}^{2}$, this is the claimed expansion. Finally, $B(\zeta_{c}^{2})<0$ and $\log(1-\zeta^{2}/\zeta_{c}^{2})\to$-$\infty$, hence $\sigma_{p}(\zeta)\to+\infty$. ∎
5.5. Continued Gram weights and edge density
The Euler identity from Lemma 5.2 continues to make sense on the slit plane once $\mathcal{G}_{p}$ has been analytically continued. This gives a canonical continuation of the scalar Gram weights.
Definition 5.11 (Continued scalar Gram weights).
For $u\in\mathbb{C}\setminus[\zeta_{c}^{2},\infty)$, define
(5.9) $\sigma_{p}^{\mathrm{cont}}(u):=\frac{1}{p}\Bigl(p+s\,u\frac{d}{du}\Bigr)^{\!2}\mathcal{G}_{p}(u).$
For $|u|<\zeta_{c}^{2}$, this agrees with the original Gram weight $\sigma_{p}(\zeta)$ under the substitution $u=\zeta^{2}$, by Lemma 5.2.
The function $\sigma_{p}^{\mathrm{cont}}$ inherits the branch cut $\zeta_{c}^{2},\infty)$ from $\mathcal{G}_{p}$. Its jump across the cut is the natural supercritical analogue of the logarithmic divergence on the subcritical side.
Definition 5.12 (Discontinuity density).
For $u>\zeta_{c}^{2}$, define
(5.10) $\rho_{p}(u):=-\frac{1}{2\pi i}\,\operatorname{Disc}\,\sigma_{p}^{\mathrm{cont}}(u)=-\frac{1}{2\pi i}\bigl(\sigma_{p}^{\mathrm{cont}}(u+i0)-\sigma_{p}^{\mathrm{cont}}(u-i0)\bigr).$
The sign convention in (5.10) is chosen so that the edge value is positive when $B(\zeta_{c}^{2})<0$, in agreement with the standard resolvent orientation.
Theorem 5.13 (Nonvanishing edge density).
The limit
$\rho_{p}(\zeta_{c}^{2}):=\lim_{u\downarrow\zeta_{c}^{2}}\rho_{p}(u)$
exists and is nonzero for all $s\geq 2$ and $p\geq 1$. More precisely,
(5.11) $\rho_{p}(\zeta_{c}^{2})=-\frac{2s^{2}}{p}\,B(\zeta_{c}^{2}).$
Proof.
Let $u>\zeta_{c}^{2}$ and set $w:=1-u/\zeta_{c}^{2}\in(-\infty,0)$. By Theorem 5.9, in a punctured neighborhood of $u=\zeta_{c}^{2}$ one has
$\mathcal{G}_{p}(u)=A(u)+B(u)\,w^{2}\log w,$
with $A$ and $B$ analytic near $\zeta_{c}^{2}$. Since $A$ and $B$ are single-valued across the cut, only the logarithm contributes to the jump. On the principal branch, $\operatorname{Disc}\,\log w=2\pi i$, $u>\zeta_{c}^{2}$, hence
(5.12) $\operatorname{Disc}\,\mathcal{G}_{p}(u)=2\pi i\,B(u)\,w^{2},\qquad u>\zeta_{c}^{2}\text{ close to }\zeta_{c}^{2}.$
Because the Euler operator has analytic coefficients away from the endpoint, it may be applied separately to the upper and lower holomorphic boundary values on the cut. Subtracting the two resulting expressions shows that it commutes with the discontinuity operation. Therefore,
$\operatorname{Disc}\,\sigma_{p}^{\mathrm{cont}}(u)=\frac{1}{p}\Bigl(p+s\,u\frac{d}{du}\Bigr)^{\!2}\operatorname{Disc}\,\mathcal{G}_{p}(u)=\frac{2\pi i}{p}\Bigl(p+s\,u\frac{d}{du}\Bigr)^{\!2}\!\bigl[B(u)w^{2}\bigr].$
Now $u=\zeta_{c}^{2}(1-w)$, so u(d/du)=(1w)(d/dw)u(d/du)=-(1-w)(d/dw). Since $B(u)=B(\zeta_{c}^{2})+O(w)$, one finds
$\Bigl(p+s\,u\frac{d}{du}\Bigr)\!\bigl[B(u)w^{2}\bigr]=-2s\,B(\zeta_{c}^{2})\,w+O(w^{2}),$
and applying the same operator once more yields
(5.13) $\Bigl(p+s\,u\frac{d}{du}\Bigr)^{\!2}\!\bigl[B(u)w^{2}\bigr]=2s^{2}\,B(\zeta_{c}^{2})+O(w),\qquad u\to\zeta_{c}^{2}.$
Substituting into (5.10) gives
$\rho_{p}(u)=-\frac{1}{p}\Bigl(2s^{2}\,B(\zeta_{c}^{2})+O(w)\Bigr),$
and letting $u\downarrow\zeta_{c}^{2}$ proves (5.11). ∎
Remark 5.14 (Matching across the analytic threshold).
Corollary 5.10 and Theorem 5.13 describe the same resonant singularity from opposite sides of the threshold. On the subcritical side $(u\uparrow\zeta_{c}^{2})$, the Euler operator converts the term
$B(u)\Bigl(1-\frac{u}{\zeta_{c}^{2}}\Bigr)^{2}\log\Bigl(1-\frac{u}{\zeta_{c}^{2}}\Bigr)$
into a bare logarithm, producing the divergence $\sigma_{p}(\zeta)\to+\infty$. On the supercritical side $(u>\zeta_{c}^{2})$, the same logarithm acquires the jump
$\operatorname{Disc}\,\log\Bigl(1-\frac{u}{\zeta_{c}^{2}}\Bigr)=2\pi i,$
and this yields the nonzero edge value $\rho_{p}(\zeta_{c}^{2})$. Thus the subcritical blow-up and the supercritical edge density are two boundary manifestations of one and the same branch-point coefficient $B(\zeta_{c}^{2})$.
To illustrate the continuation across the critical point, we plot in Figure 3 the continued scalar Gram weight $\sigma_{p}^{\mathrm{cont}}(u)$, which agrees with the subcritical Gram weight for $u<\zeta_{c}^{2}$, together with the supercritical discontinuity density $\rho_{p}(u)$ for several values of $p$. The lower panels show that the edge value
$\rho_{p}(\zeta_{c}^{2})=-\frac{2s^{2}}{p}\,B(\zeta_{c}^{2})$
is positive, in agreement with the local expansion at the branch point. At the same time, the additional curves with larger $p$ show that $\rho_{p}(u)$ need not remain positive for all $u>\zeta_{c}^{2}$: after starting from a positive edge value, the density may cross zero and become negative further along the supercritical branch. Thus the sign of $\rho_{p}$ is controlled locally near $u=\zeta_{c}^{2}$, but is not globally fixed on the whole continuation domain.
Figure 3. Continued Gram weight $\sigma_{p}^{\mathrm{cont}}(u)$ and discontinuity density $\rho_{p}(u)$ for $s=3,5$. Top row: the analytically continued Gram weight $\sigma_{p}^{\mathrm{cont}}(u)$ on the subcritical side $0<u<\zeta_{c}^{2}$. Bottom row: the discontinuity density $\rho_{p}(u)$ on the supercritical side $u>\zeta_{c}^{2}$. The open circles at $u=\zeta_{c}^{2}$ mark the edge values $\rho_{p}(\zeta_{c}^{2})>0$. For larger values of $p$, the density becomes negative away from the edge, showing that positivity at $u=\zeta_{c}^{2}$ does not persist on the entire supercritical branch.
This numerical behavior is consistent with the edge asymptotics proved above, while showing that the positivity statement is local in $u$ and does not extend to the full supercritical branch.
5.6. Cauchy–Stieltjes representation
The hypergeometric continuation of $\mathcal{G}_{p}$ on the slit plane produces canonical boundary values across the cut and therefore a Cauchy–Stieltjes representation. This representation does not require any positivity assumption and is valid for every $p\geq 1$.
Lemma 5.15 (Growth at infinity).
Fix $p\geq 1$. As $|u|\to\infty$ in $\mathbb{C}\setminus[\zeta_{c}^{2},\infty)$,
(5.14) $\mathcal{G}_{p}(u)=O\bigl(|u|^{-p/s}(1+\log|u|)\bigr).$
Consequently, for
$F_{p}(\eta):=\eta^{-1}\mathcal{G}_{p}(1/\eta),$
one has
(5.15) $F_{p}(\eta)=O\bigl(|\eta|^{p/s-1}(1+\log(1/|\eta|))\bigr),\qquad\eta\to 0,\ \eta\notin[0,1/\zeta_{c}^{2}].$
Proof.
By Proposition 5.4, $\mathcal{G}_{p}(u)$ is a generalized hypergeometric function in the variable $\xi=u/\zeta_{c}^{2}$, with numerator parameters
$\alpha_{k}=\frac{p+k}{s},\qquad k=0,\dots,s-1,$
each occurring with multiplicity $2$. The generalized hypergeometric differential equation has regular singular points at $\xi=0,1,\infty$, and the local exponents at $\xi=\infty$ are precisely the numerator parameters; see [18, §16.11]. Because each exponent appears with multiplicity $2$, the asymptotic expansion contains at most one logarithm per exponent:
$\mathcal{G}_{p}(u)=\sum_{k=0}^{s-1}u^{-\alpha_{k}}\bigl(A_{k}\log u+B_{k}\bigr)+O\!\bigl(|u|^{-(p+1)/s}\log|u|\bigr),$
uniformly on every closed subsector of the slit plane. The smallest exponent is $\alpha_{0}=p/s$, which yields (5.14). Substituting $u=1/\eta$ gives (5.15). ∎
Proposition 5.16 (Stieltjes representation).
There exists a finite real signed Borel measure $\nu_{p}$ of bounded total variation, supported on $0,1/\zeta_{c}^{2}]$, such that
(5.16) $\mathcal{G}_{p}(u)=\int_{0}^{1/\zeta_{c}^{2}}\frac{d\nu_{p}(t)}{1-ut},\qquad u\in\mathbb{C}\setminus[\zeta_{c}^{2},\infty).$
Moreover,
$\nu_{p}\bigl([0,1/\zeta_{c}^{2}]\bigr)=\mathcal{G}_{p}(0)=1,$
and $\nu_{p}$ is absolutely continuous on the open interval $(0,1/\zeta_{c}^{2})$, with density
(5.17) $\frac{d\nu_{p}}{dt}(t)=\frac{1}{2\pi i\,t}\,\operatorname{Disc}\,\mathcal{G}_{p}\Bigl(\frac{1}{t}\Bigr),\qquad 0<t<1/\zeta_{c}^{2}.$
Proof.
Set
$F_{p}(\eta):=\eta^{-1}\mathcal{G}_{p}(1/\eta).$
Since $\mathcal{G}_{p}$ is holomorphic on $\mathbb{C}\setminus[\zeta_{c}^{2},\infty)$, the function $F_{p}$ is holomorphic on $\mathbb{C}\setminus[0,1/\zeta_{c}^{2}]$. From the Taylor expansion $\mathcal{G}_{p}(u)=1+O(u)$ at $u=0$, one obtains
$F_{p}(\eta)=\eta^{-1}+O(\eta^{-2}),\qquad|\eta|\to\infty.$
On the other hand, Lemma 5.15 gives
$F_{p}(\eta)=o(|\eta|^{-1}),\qquad\eta\to 0,$
away from the cut.
Let $\eta\in\mathbb{C}\setminus[0,1/\zeta_{c}^{2}]$, and apply Cauchy’s theorem to a keyhole contour around the interval $0,1/\zeta_{c}^{2}]$, with outer radius $R$ and inner radius $\varepsilon$. The contribution of the outer circle vanishes as $R\to\infty$ because $F_{p}(\xi)=O(\xi^{-1})$, while the inner-circle contribution vanishes as $\varepsilon\to 0$ by the bound (5.15). Passing to the limit yields
(5.18) $F_{p}(\eta)=\frac{1}{2\pi i}\int_{0}^{1/\zeta_{c}^{2}}\frac{\operatorname{Disc}\,F_{p}(t)}{t-\eta}\,dt.$
Define a signed Borel measure on $0,1/\zeta_{c}^{2}]$ by
$d\nu_{p}(t):=\frac{1}{2\pi i}\,\operatorname{Disc}\,F_{p}(t)\,dt.$
Its total variation is finite, and this is exactly where the endpoint estimates enter. If $0<t<1/\zeta_{c}^{2}$ is close to the right endpoint and $u=1/t$, then $u>\zeta_{c}^{2}$ and Theorem 5.9 gives
$\operatorname{Disc}\,\mathcal{G}_{p}(u)=2\pi i\,B(u)\Bigl(1-\frac{u}{\zeta_{c}^{2}}\Bigr)^{2}.$
Hence
$\operatorname{Disc}\,F_{p}(t)=t^{-1}\operatorname{Disc}\,\mathcal{G}_{p}(1/t)=O\Bigl(\Bigl(\frac{1}{\zeta_{c}^{2}}-t\Bigr)^{2}\Bigr),\qquad t\uparrow\frac{1}{\zeta_{c}^{2}},$
so the density is integrable at the right endpoint. Near $t=0$, (5.15) gives
$\operatorname{Disc}\,F_{p}(t)=O\bigl(t^{p/s-1}(1+|\log t|)\bigr),$
which is integrable because $p/s>0$. Therefore $|\operatorname{Disc}F_{p}(t)|$ is integrable on $0,1/\zeta_{c}^{2}]$, and the resulting measure $\nu_{p}$ has bounded total variation. Thus (5.18) becomes
$F_{p}(\eta)=\int_{0}^{1/\zeta_{c}^{2}}\frac{d\nu_{p}(t)}{t-\eta}.$
Substituting $\eta=1/u$ and multiplying by $u$ gives (5.16).
Next,
$\operatorname{Disc}\,F_{p}(t)=t^{-1}\operatorname{Disc}\,\mathcal{G}_{p}(1/t),$
which yields the jump formula (5.17). Evaluating (5.16) at $u=0$ gives
$\nu_{p}\bigl([0,1/\zeta_{c}^{2}]\bigr)=\mathcal{G}_{p}(0)=1.$
Finally, $\nu_{p}$ is real because $\mathcal{G}_{p}$ has real Taylor coefficients and therefore satisfies Schwarz reflection on the slit plane. ∎
Remark 5.17.
For general $p$, the representation (5.16)–(5.17) is used only as an analytic description of the continued scalar Gram data. Positivity is not required at this stage. The operator-theoretic interpretation enters only later, in the range $1\leq p\leq s$, where $\nu_{p}$ becomes a positive measure and one can pass to the Jacobi/OPRL framework.
5.7. Jacobi realization in the positive range
For general $p$, Proposition 5.16 yields only a signed representing measure for $\mathcal{G}_{p}$. From this point onward we restrict to the positive-measure range $1\leq p\leq s$. In that range, positivity is supplied by an external Hausdorff-moment theorem for Raney numbers. We record exactly the part of that input used here, in the normalization relevant for the present paper, before passing to the Jacobi/OPRL framework.
Lemma 5.18 (External Hausdorff-moment input in the present normalization).
Assume $s\geq 2$ and $0\leq p\leq s$. Then there exists a probability measure $\nu_{s,p}$ supported on
$0,\tau_{s}],\qquad\tau_{s}:=\frac{s^{s}}{(s-1)^{s-1}}=\zeta_{c}^{-1},$
such that
$R_{s,p}(n)=\int_{0}^{\tau_{s}}x^{n}\,d\nu_{s,p}(x),\qquad n\geq 0.$
Equivalently, the rescaled sequence $\{R_{s,p}(n)\zeta_{c}^{n}\}_{n\geq 0}$ is a Hausdorff moment sequence on $[0,1]$.
This lemma is the only place where the external Raney-moment positivity input is used. In what follows, we use only the existence of the positive measure and the support endpoint $\tau_{s}=\zeta_{c}^{-1}$; no explicit formula for the density is needed. These facts are supplied by [14, Thm. 5] together with [15]. For background on the corresponding Raney distributions, see also [7].
Proof.
By Proposition 2.4,
$R_{s,p}(n)=\frac{p}{sn+p}\binom{sn+p}{n}$.
In the notation of [14], this is exactly the Raney sequence $A_{n}(s,p)$. Therefore [14, Thm. 5], as corrected in [15], applies with the parameter pair $(p,r)=(s,p)$: the hypotheses are satisfied because $s\geq 1$ and $0\leq p\leq s$. It follows that $\{R_{s,p}(n)\}_{n\geq 0}$ is represented by a probability measure supported on $0,\tau_{s}]$, with
$\tau_{s}=\frac{s^{s}}{(s-1)^{s-1}}.$
In the normalization of the present paper, $\tau_{s}=\zeta_{c}^{-1}$, which gives exactly the stated moment representation. After the rescaling $x=\tau_{s}y$, this is equivalent to the Hausdorff moment statement for $\{R_{s,p}(n)\zeta_{c}^{n}\}_{n\geq 0}$ on $[0,1]$. ∎
Proposition 5.19 (Positivity for $1\leq p\leq s$).
Assume $1\leq p\leq s$. Then the measure $\nu_{p}$ in Proposition 5.16 can be chosen to be a positive probability measure supported on $0,1/\zeta_{c}^{2}]$. Equivalently, after the rescaling $t\mapsto t\,\zeta_{c}^{2}$, the sequence
$\{R_{s,p}(n)^{2}\,\zeta_{c}^{2n}\}_{n\geq 0}$
is a Hausdorff moment sequence.
Proof.
By Lemma 5.18, for $1\leq p\leq s$ there exists a probability measure $\nu_{s,p}$ supported on $0,\tau_{s}]=[0,\zeta_{c}^{-1}]$ such that
$R_{s,p}(n)=\int_{0}^{\tau_{s}}x^{n}\,d\nu_{s,p}(x),\qquad n\geq 0.$
Let $\widetilde{\nu}_{p}$ be the pushforward of $\nu_{s,p}\otimes\nu_{s,p}$ under the multiplication map
$(x,y)\mapsto xy.$
Then $\widetilde{\nu}_{p}$ is a probability measure supported on $0,\tau_{s}^{2}]=[0,\zeta_{c}^{-2}]$, and for every $n\geq 0$,
$\int_{0}^{1/\zeta_{c}^{2}}t^{n}\,d\widetilde{\nu}_{p}(t)=\iint(xy)^{n}\,d\nu_{s,p}(x)\,d\nu_{s,p}(y)=\biggl(\int_{0}^{\tau_{s}}x^{n}\,d\nu_{s,p}(x)\biggr)^{2}=R_{s,p}(n)^{2}.$
Thus the rescaled sequence $\{R_{s,p}(n)^{2}\zeta_{c}^{2n}\}_{n\geq 0}$ is a Hausdorff moment sequence on $[0,1]$. Consequently,
$\int_{0}^{1/\zeta_{c}^{2}}\frac{d\widetilde{\nu}_{p}(t)}{1-ut}=\sum_{n\geq 0}R_{s,p}(n)^{2}\,u^{n}=\mathcal{G}_{p}(u),\qquad|u|<\zeta_{c}^{2}.$
Since both sides are holomorphic on $\mathbb{C}\setminus[\zeta_{c}^{2},\infty)$, the identity theorem extends this representation to the whole slit plane.
On the other hand, Proposition 5.16 already provides a finite signed measure $\nu_{p}$ on $0,1/\zeta_{c}^{2}]$ with the same Stieltjes transform:
$\mathcal{G}_{p}(u)=\int_{0}^{1/\zeta_{c}^{2}}\frac{d\nu_{p}(t)}{1-ut},\qquad u\in\mathbb{C}\setminus[\zeta_{c}^{2},\infty).$
Hence
$\int_{0}^{1/\zeta_{c}^{2}}\frac{d(\nu_{p}-\widetilde{\nu}_{p})(t)}{1-ut}=0$
on the slit plane. By uniqueness of the Cauchy–Stieltjes transform of a finite compactly supported measure, one has $\nu_{p}=\widetilde{\nu}_{p}$. Therefore the measure in Proposition 5.16 is in fact positive, and being a probability measure it has total mass $1$. ∎
Once positivity is available, the standard orthogonal-polynomial machinery produces a Jacobi operator with spectral measure $\nu_{p}$.
Definition 5.20 (Jacobi operator associated with $\mathcal{G}_{p}$).
Assume $1\leq p\leq s$, and let $\{P_{n}^{(p)}\}_{n\geq 0}$ be the monic orthogonal polynomials with respect to the positive measure $\nu_{p}$ from Proposition 5.19. Then there exist coefficients
$a_{n}^{(p)}>0,\qquad b_{n}^{(p)}\in\mathbb{R},$
such that
(5.19) $tP_{n}^{(p)}(t)=P_{n+1}^{(p)}(t)+b_{n}^{(p)}P_{n}^{(p)}(t)+(a_{n}^{(p)})^{2}P_{n-1}^{(p)}(t),$
with $P_{-1}^{(p)}\equiv 0$. The associated Jacobi operator $J_{p}$ is the bounded self-adjoint operator on $\ell^{2}(\mathbb{N}_{0})$ with tridiagonal matrix
$J_{p}=\begin{pmatrix}b_{0}^{(p)}&a_{1}^{(p)}&0&\cdots\\ a_{1}^{(p)}&b_{1}^{(p)}&a_{2}^{(p)}&\ddots\\ 0&a_{2}^{(p)}&b_{2}^{(p)}&\ddots\\ \vdots&\ddots&\ddots&\ddots\end{pmatrix}$.
Since $\textrm{supp}(\nu_{p})\subset[0,1/\zeta_{c}^{2}]$, one has
$\sigma(J_{p})\subset[0,1/\zeta_{c}^{2}].$
Theorem 5.21 (Weyl function identity).
Assume $1\leq p\leq s$. Then $\mathcal{G}_{p}$ is the Weyl $m$-function of $J_{p}$ in the reciprocal spectral parameter:
(5.20) $\mathcal{G}_{p}(u)=\langle e_{0},(I-uJ_{p})^{-1}e_{0}\rangle,\qquad u\in\mathbb{C}\setminus[\zeta_{c}^{2},\infty),$
where $e_{0}=(1,0,0,\dots)^{T}$. Equivalently, for $u\neq 0$, the formula holds whenever $u^{-1}\in\rho(J_{p})$.
Proof.
By the spectral theorem,
$\langle e_{0},(I-uJ_{p})^{-1}e_{0}\rangle=\int_{0}^{1/\zeta_{c}^{2}}\frac{d\nu_{p}(t)}{1-ut},\qquad u\in\mathbb{C}\setminus[\zeta_{c}^{2},\infty),$
where $\nu_{p}$ is the spectral measure of $J_{p}$ at $e_{0}$. By Proposition 5.19, this is precisely the positive representing measure of $\mathcal{G}_{p}$, and the right-hand side equals $\mathcal{G}_{p}(u)$ by (5.16). This proves (5.20); see, for example, [1, Ch. III]. ∎
Theorem 5.22 (Stieltjes–Perron inversion).
Assume $1\leq p\leq s$. Then the positive measure $\nu_{p}$ from Proposition 5.19 is absolutely continuous on $(0,1/\zeta_{c}^{2})$, and its density is
(5.21) $\varrho_{p}(t)=\frac{1}{\pi t}\,\Im\,\mathcal{G}_{p}\Bigl(\frac{1}{t}+i0\Bigr)=\frac{1}{2\pi i\,t}\,\operatorname{Disc}\,\mathcal{G}_{p}\Bigl(\frac{1}{t}\Bigr),\qquad 0<t<1/\zeta_{c}^{2}.$
Proof.
Proposition 5.16 constructs a representing measure $\nu_{p}$ that is absolutely continuous on the open interval $(0,1/\zeta_{c}^{2})$, with density given by the jump formula (5.17). In the range $1\leq p\leq s$, Proposition 5.19 shows that the same function $\mathcal{G}_{p}$ also admits a positive representing measure. By uniqueness of the Stieltjes transform of a finite compactly supported measure, this positive measure coincides with the $\nu_{p}$ from Proposition 5.16. Hence $\nu_{p}$ is positive and its density in the spectral variable $t$ can therefore be written in the Stieltjes–Perron form displayed in (5.21); see, for example, [1, Ch. III]. ∎
Remark 5.23 (Nodal structure of the soft eigenvectors).
The Jacobi realization suggests a further structural property of the compact remainder $\widetilde{C}^{(q)}_{*}$ introduced in Theorem 4.6. Let $\phi_{2}^{(q)},\phi_{3}^{(q)},\ldots$ denote the eigenvectors of $\widetilde{C}^{(q)}_{*,\perp}$, ordered by decreasing eigenvalue $\mu_{2,*}^{(q)}\geq\mu_{3,*}^{(q)}\geq\cdots$, and view each $\phi_{k}^{(q)}$ as a function on the lattice $\{p_{j}=q+js\}_{j\geq 0}$. Numerically, $\phi_{k}^{(q)}$ exhibits exactly k1k-1 sign changes in the index $j$ (Figure 4), a pattern stable across all values of $s$, $q$, and $\beta$ tested. This oscillation count is the hallmark of classical Sturm oscillation for self-adjoint Jacobi matrices [1, Ch. III]. For the range $1\leq p\leq s$, each scalar factor $\mathcal{G}_{p_{j}}$ admits such a Jacobi realization by Theorem 5.21. The observed nodal property of the full matrix $\widetilde{C}^{(q)}_{*}$ is therefore consistent with, though not a direct corollary of, the scalar Jacobi framework, since the Gram entries mix contributions from different Jacobi operators $J_{p_{j}}$. We leave a rigorous nodal theorem for $\widetilde{C}^{(q)}_{*}$ as an open problem.
Figure 4. Structure of the first soft modes of $\widetilde{C}^{(q)}_{*,\perp}$ for $s=3,5$, with $q=1$, $\beta=1$, and $N=40$. The curves show the signed components of the eigenvectors $\phi_{2}^{(q)},\phi_{3}^{(q)},\phi_{4}^{(q)},\phi_{5}^{(q)}$, corresponding to the soft eigenvalues $\mu_{2,*}^{(q)},\mu_{3,*}^{(q)},\mu_{4,*}^{(q)},\mu_{5,*}^{(q)}$. As the mode index increases, the eigenvectors become more oscillatory while remaining concentrated near low values of the lattice index $p_{j}=q+js$.
The results of this section provide the analytic framework for the intermediate regime $\zeta_{c}<\zeta<\zeta_{\mathrm{univ}}$. The generating functions $\mathcal{G}_{p}(u)$ extend to the slit plane via their hypergeometric structure, have a resonant singularity of parametric excess $\gamma=2$ at the branch point, and admit a nonvanishing edge density; for $1\leq p\leq s$, $\mathcal{G}_{p}$ also has a Jacobi–Weyl realization. Thus the scalar spectral data persist beyond the loss of the weighted compact operator picture. Geometric singularity on the domain boundary occurs only at $\zeta=\zeta_{\mathrm{univ}}$: for $s\geq 3$ this is cusp formation, whereas for $s=2$ the critical map degenerates to the classical Joukowski segment. We treat this geometric threshold next in Section 6.
6. The geometric threshold $\zeta_{\mathrm{univ}}$
We now return to the geometric side of the problem. For the polynomial map
$f(w)=rw+aw^{1-s},\qquad\zeta:=\frac{a}{r}>0,$
the parameter
$\zeta_{\mathrm{univ}}=\frac{1}{s-1}$
marks the loss of univalence of the exterior conformal map. This threshold is independent of the weighted spectral theory developed in Section 4 and of the scalar continuation theory developed in Section 5. The purpose of the present section is to compare these two notions of criticality.
We first characterize the geometric transition itself. Below $\zeta_{\mathrm{univ}}$ the boundary is a smooth Jordan curve. At $\zeta_{\mathrm{univ}}$, the critical geometry depends on $s$: for $s\geq 3$ the boundary develops semicubical cusps, whereas for $s=2$ the Joukowski trace degenerates to a line segment. Above $\zeta_{\mathrm{univ}}$, the map is no longer injective; for $s\geq 3$ the boundary self-intersects, while for $s=2$ the unit-circle trace remains an ellipse although the exterior map ceases to be one-to-one. We then show that
$\zeta_{c}<\zeta_{\mathrm{univ}},$
so the logarithmic spectral instability from Theorems 4.6 and 4.7 occurs while the conformal map is still univalent. Finally, we prove that the analytically continued scalar Gram data remain finite at $u=\zeta_{\mathrm{univ}}^{2}$. Thus the spectral singularity at $\zeta_{c}$ and the geometric singularity at $\zeta_{\mathrm{univ}}$ are genuinely distinct.
6.1. Geometric threshold
For the comparison with the analytic threshold $\zeta_{c}$, we only need the location of the geometric threshold at which the exterior map ceases to be univalent. The finer local description of the boundary singularity at the threshold is classical and will not be used below.
Proposition 6.1 (Geometric threshold).
Let
$f(w)=rw+aw^{1-s}=r\bigl(w+\zeta w^{1-s}\bigr),\qquad s\geq 2,\qquad\zeta:=a/r>0.$
Then the univalence threshold on the exterior disk is
$\zeta_{\mathrm{univ}}=\frac{1}{s-1}.$
More precisely:
  1. (1)
    if $\zeta<\zeta_{\mathrm{univ}}$, then $f$ is univalent on $\{|w|>1\}$ and extends injectively to $|w|=1$;
  2. (2)
    if $\zeta>\zeta_{\mathrm{univ}}$, then $f$ is not univalent on $\{|w|>1\}$;
  3. (3)
    if $\zeta=\zeta_{\mathrm{univ}}$, then
    $f^{\prime}(w)=r\bigl(1-(s-1)\zeta w^{-s}\bigr)$
    vanishes exactly at the $s$ points $w^{s}=1$ on the unit circle.
Proof.
For $w=e^{i\theta}$, write
$z(\theta):=f(e^{i\theta})=r\bigl(e^{i\theta}+\zeta e^{-i(s-1)\theta}\bigr).$
Suppose $z(\theta)=z(\phi)$. Then $e^{i\theta}-e^{i\phi}=\zeta(e^{-i(s-1)\phi}-e^{-i(s-1)\theta})$. Taking absolute values and setting $\delta=(\theta-\phi)/2$, we obtain
$|\sin\delta|=\zeta\,|\sin((s-1)\delta)|\leq(s-1)\zeta\,|\sin\delta|.$
Hence, if $(s-1)\zeta<1$, then $\sin\delta=0$, so $\theta\equiv\phi\pmod{2\pi}$. Thus the boundary trace is injective. Moreover, $f^{\prime}(w)=r\bigl(1-(s-1)\zeta w^{-s}\bigr)$, so for $(s-1)\zeta<1$ all critical points satisfy
$|w|^{s}=(s-1)\zeta<1.$
Therefore $f^{\prime}(w)\neq 0$ on $|w|\geq 1$, so $f$ is a holomorphic local biholomorphism on $\{|w|>1\}$. Since $f(w)=rw+O(1)$ as $w\to\infty$, the map is proper as a map from the exterior disk to its image. Hence $f$ is a covering map of finite degree onto its image. The normalization at infinity shows that this degree is $1$: for $|z|$ sufficiently large, the equation $f(w)=z$ has exactly one solution in $|w|>1$, namely the branch with $w=z/r+O(1)$. Therefore the covering is trivial, so $f$ is univalent on $\{|w|>1\}$ and extends injectively to $|w|=1$.
If $(s-1)\zeta>1$, then the equation $f^{\prime}(w)=0$ has solutions with
$|w|^{s}=(s-1)\zeta>1,$
so $f^{\prime}$ vanishes inside the exterior domain. Since a holomorphic injective map cannot have a critical point, $f$ is not univalent on $\{|w|>1\}$.
Finally, if $(s-1)\zeta=1$, then $f^{\prime}(w)=0$ is equivalent to $w^{s}=1$, so the critical points lie exactly on the unit circle. ∎
Remark 6.2.
At the critical value $\zeta=\zeta_{\mathrm{univ}}$, the boundary singularity is standard. For $s\geq 3$, the image of $|w|=1$ develops $s$ semicubical cusps; for $s=2$, the Joukowski trace degenerates to a line segment. These local descriptions are classical and are not needed for the spectral analysis below.
6.2. Separation of the analytic and geometric thresholds
With the geometric threshold identified, we now compare it with the analytic threshold $\zeta_{c}=(s-1)^{s-1}/s^{s}$.
Proposition 6.3 (Separation of thresholds).
For every $s\geq 2$,
$\zeta_{c}<\zeta_{\mathrm{univ}}.$
Equivalently,
(6.1) $\frac{\zeta_{c}}{\zeta_{\mathrm{univ}}}=\Bigl(\frac{s-1}{s}\Bigr)^{s}<1.$
In particular, the logarithmic spectral transition described by Theorems 4.6 and 4.7 takes place while the conformal map is still univalent and the boundary is still smooth.
Proof.
The identity (6.1) is immediate:
$\frac{\zeta_{c}}{\zeta_{\mathrm{univ}}}=(s-1)\frac{(s-1)^{s-1}}{s^{s}}=\Bigl(\frac{s-1}{s}\Bigr)^{s}.$
Since (s1)/s<1(s-1)/s<1, the right-hand side is strictly less than 11. ∎
6.3. Regularity of the continued scalar data at $\zeta_{\mathrm{univ}}$
We finally show that the continued scalar data do not develop any new singularity at the geometric threshold. This is the final step in separating analytic and geometric criticality.
Since $\zeta_{\mathrm{univ}}>\zeta_{c}$, the point
$u_{\mathrm{univ}}:=\zeta_{\mathrm{univ}}^{2}$
lies on the branch cut $\zeta_{c}^{2},\infty)$ of the analytically continued functions $\mathcal{G}_{p}(u)$. However, $u_{\mathrm{univ}}$ is an interior point of the cut, not its endpoint. The only singular endpoint produced by the hypergeometric continuation is $u=\zeta_{c}^{2}$. Thus one expects finite lateral values at $u=u_{\mathrm{univ}}$, and the next proposition confirms this. The key point is that every interior point of the cut is an ordinary point of the hypergeometric differential equation.
Lemma 6.4 (Interior points of the cut are ordinary points).
Fix $p\geq 1$ and $u_{0}>\zeta_{c}^{2}$. Let $\mathcal{G}_{p,+}$ and $\mathcal{G}_{p,-}$ denote the analytic continuations of the germ of $\mathcal{G}_{p}$ from $u=0$ to a neighborhood of $u_{0}$ through the upper and lower half-planes, respectively. Then each branch extends holomorphically to a full disk centered at $u_{0}$. In particular,
$\mathcal{G}_{p,\pm}(u_{0}),\qquad\mathcal{G}^{\prime}_{p,\pm}(u_{0}),\qquad\mathcal{G}^{\prime\prime}_{p,\pm}(u_{0})$
are finite.
Proof.
By Proposition 5.4, $\mathcal{G}_{p}$ satisfies the generalized hypergeometric differential equation in the variable $\xi=\zeta_{c}^{-2}u$. This equation is Fuchsian, and its only finite singular points are $\xi=0$ and $\xi=1$; equivalently, in the variable $u$ the only finite singular points are $u=0$ and $u=\zeta_{c}^{2}$; see [18, §16.8,§16.11]. Hence every $u_{0}>\zeta_{c}^{2}$ with $u_{0}\neq\zeta_{c}^{2}$ is an ordinary point of the differential equation.
Choose $r>0$ so small that the closed disk $\overline{D(u_{0},r)}$ avoids $\{0,\zeta_{c}^{2}\}$. On $D(u_{0},r)$, the differential equation has analytic coefficients. Standard local ODE theory therefore implies that any solution defined on a connected open subset of $D(u_{0},r)$ extends uniquely to a holomorphic solution on the whole disk. Applying this to the branches $\mathcal{G}_{p,+}$ and $\mathcal{G}_{p,-}$, initially defined on the upper and lower half-disks, gives the claimed holomorphic extensions. Their values and first two derivatives at $u_{0}$ are therefore finite. ∎
Proposition 6.5 (Regularity at the geometric threshold).
For every $p\geq 1$, the lateral values
$\mathcal{G}_{p}(u\pm i0),\qquad\sigma_{p}^{\mathrm{cont}}(u\pm i0),$
are finite for every $u>\zeta_{c}^{2}$. In particular, they are finite at
$u=\zeta_{\mathrm{univ}}^{2},$
and the discontinuity density satisfies
$\rho_{p}(\zeta_{\mathrm{univ}}^{2})<\infty.$
Proof.
Fix $u_{0}>\zeta_{c}^{2}$. By Lemma 6.4, each lateral branch of $\mathcal{G}_{p}$ extends holomorphically to a neighborhood of $u_{0}$. Hence the limits
$\mathcal{G}_{p}(u_{0}\pm i0)$
exist and are finite, and so do the first two lateral derivatives.
Now $\sigma_{p}^{\mathrm{cont}}$ is obtained from $\mathcal{G}_{p}$ by the Euler operator
$\sigma_{p}^{\mathrm{cont}}(u)=\frac{1}{p}\Bigl(p+s\,u\frac{d}{du}\Bigr)^{2}\mathcal{G}_{p}(u),$
whose coefficients are analytic at every finite $u$. Therefore $\sigma_{p}^{\mathrm{cont}}(u_{0}\pm i0)$ is finite for every $u_{0}>\zeta_{c}^{2}$, and in particular at $u_{0}=\zeta_{\mathrm{univ}}^{2}$.
Finally, Definition 5.12 gives
$\rho_{p}(u_{0})=-\frac{1}{2\pi i}\bigl(\sigma_{p}^{\mathrm{cont}}(u_{0}+i0)-\sigma_{p}^{\mathrm{cont}}(u_{0}-i0)\bigr),$
which is therefore finite as well.∎
The three regimes of the problem may now be summarized as follows: (i) For $0<\zeta<\zeta_{c}$, each symmetry sector is described by a compact weighted Gram operator with one stiff eigenvalue diverging logarithmically as $\zeta\uparrow\zeta_{c}$, while the remaining sectorial spectrum stays bounded. Equivalently, there are at most $s$ logarithmically diverging eigenvalues globally. (ii) For $\zeta_{c}<\zeta<\zeta_{\mathrm{univ}}$, the weighted operator realization breaks down, but the scalar data $\mathcal{G}_{p}$, $\sigma_{p}^{\mathrm{cont}}$, and $\rho_{p}$ admit a canonical analytic continuation, with Jacobi–Weyl interpretation for $1\leq p\leq s$. (iii) At $\zeta=\zeta_{\mathrm{univ}}$, the conformal map loses univalence; for $s\geq 3$ the boundary develops cusp singularities, whereas for $s=2$ the critical map degenerates to the Joukowski segment. In both cases, the continued scalar data remain finite.
This completes the proof that analytic criticality and geometric criticality are distinct in the symmetric one-harmonic family.
Conclusion
We have shown that for the $s$-fold symmetric one-harmonic polynomial family the first spectral instability of the mixed Hessian of the dispersionless Toda $\tau$-function is governed by analytic criticality of the inverse map, not by the later geometric loss of univalence of the conformal map. The square-root singularity of the inverse branch produces the borderline coefficient decay that drives the transition.
On the subcritical side, for any fixed $\beta>0$, the weighted Hilbert realization of Definition 4.1 separates this mechanism into one stiff direction and a bounded soft sector in each symmetry block. Accordingly, in each such weighted realization each sector contains exactly one logarithmically diverging eigenvalue, while the remaining sectorial spectrum stays bounded; after removing the spike direction, each fixed soft eigenvalue converges to the corresponding eigenvalue of a compact limiting remainder. The intrinsic coefficient-level content is the existence of a single dominant singular direction in each sector from which these weighted spikes are realized. Beyond the analytic threshold, the weighted compact realization ceases to apply, but the scalar Gram data remain well defined after analytic continuation. They are described by generalized hypergeometric functions, admit a Cauchy–Stieltjes representation, and in the positive range fit naturally into the Jacobi/Weyl framework.
The geometric threshold lies strictly beyond the analytic one. Consequently, the Toda Hessian becomes spectrally singular while the conformal map is still univalent and the boundary is still smooth, and the continued scalar data have finite upper and lower lateral boundary values at the univalence threshold. In this sense, analytic and geometric criticality are genuinely distinct in the present family.
The results also admit a natural heuristic interpretation in adjacent frameworks. In the Laplacian-growth language, the rank-one logarithmic spike singles out a distinguished stiff deformation mode before any geometric singularity appears. In the matrix-model or potential-theoretic language, the proved statement is that the mixed second-order response concentrates onto a single dominant sectorial direction at criticality. These interpretive remarks are not used in the proofs and are included only to indicate the structural scope of the result.
The argument isolates a structural mechanism that should persist beyond the one-harmonic leaf: a dominant branch point, borderline coefficient decay, and positive Gram structure. Several extensions therefore suggest themselves. For more general polynomial conformal maps one expects several competing dominant orbits and hence a finite-rank version of the present instability mechanism. It would also be natural to understand the bounded soft sector more conceptually, possibly through hidden Jacobi or total-positivity structure, and to sharpen the leading-order theory by deriving asymptotics for spectral gaps, sector dependence, and soft eigenvalues.
Acknowledgments
The study was implemented in the framework of the Basic Research Program at HSE University in 2026.
Appendix A Asymptotics of Raney numbers
This appendix provides explicit Stirling–type asymptotics for the Raney numbers in (2.6). These formulas are used in the Gram estimates and in Appendix D.
Proposition A.1 (Explicit $m^{-3/2}$ asymptotics).
Fix integers $s\geq 2$, $p\geq 1$, and set
$\zeta_{c}:=\frac{(s-1)^{s-1}}{s^{s}},\qquad M:=\frac{s}{s-1}.$
As $m\to\infty$,
(A.1) $R_{s,p}(m)=A_{s,p}\,\zeta_{c}^{-m}\,m^{-3/2}\bigl(1+O(m^{-1})\bigr),\qquad A_{s,p}=\frac{p}{\sqrt{2\pi}}\,\frac{s^{\,p-\frac{1}{2}}}{(s-1)^{\,p+\frac{1}{2}}}=\frac{p\,M^{p}}{\sqrt{2\pi s(s-1)}}.$
Consequently,
(A.2) $R_{s,p}(m)^{2}=\frac{p^{2}}{2\pi}\,\frac{s^{\,2p-1}}{(s-1)^{\,2p+1}}\,\zeta_{c}^{-2m}\,m^{-3}\bigl(1+O(m^{-1})\bigr).$
Proof.
Starting from (2.6) and writing factorials via Gamma functions,
$R_{s,p}(m)=\frac{p}{sm+p}\,\frac{\Gamma(sm+p+1)}{\Gamma(m+1)\,\Gamma((s-1)m+p+1)}=\frac{p\,\Gamma(sm+p)}{\Gamma(m+1)\,\Gamma((s-1)m+p+1)}.$
For fixed $s,p$ we apply Stirling’s formula
$\Gamma(z)=\sqrt{2\pi}\,z^{z-\frac{1}{2}}e^{-z}\bigl(1+O(z^{-1})\bigr),\qquad z\to+\infty,$
to the three Gamma factors with arguments $sm+p$, $m+1$, and (s1)m+p+1(s-1)m+p+1. The exponential terms $e^{-sm}e^{m}e^{(s-1)m}$ cancel, while the power terms produce
$\frac{(sm)^{sm+p-\frac{1}{2}}}{m^{m+\frac{1}{2}}((s-1)m)^{(s-1)m+p+\frac{1}{2}}}=\Bigl(\frac{s^{s}}{(s-1)^{s-1}}\Bigr)^{m}m^{-3/2}\frac{s^{p-\frac{1}{2}}}{(s-1)^{p+\frac{1}{2}}}\bigl(1+O(m^{-1})\bigr).$
Since $\zeta_{c}^{-1}=s^{s}/(s-1)^{s-1}$, this gives (A.1); squaring yields (A.2). ∎
Proposition A.2 (Uniform upper bound).
There exists a constant $C_{s}>0$ such that for all integers $p\geq 1$ and $m\geq 1$,
(A.3) $R_{s,p}(m)\leq C_{s}\,p\,M^{p}\,\zeta_{c}^{-m}\,m^{-3/2}.$
Moreover, the same bound holds with $m^{-3/2}$ replaced by $(1+m)^{-3/2}$ for all $m\geq 0$.
Proof.
We start from the closed form (2.6):
$R_{s,p}(m)=\frac{p}{sm+p}\binom{sm+p}{m},\qquad m\geq 1.$
Set $N:=sm+p$ and K:=Nm=(s1)m+pK:=N-m=(s-1)m+p. Using a standard two-sided Stirling bound (e.g. Robbins’ bound), there is an absolute constant $C>0$ such that
$\binom{N}{m}\leq C\,\frac{N^{N+\frac{1}{2}}}{m^{m+\frac{1}{2}}\,K^{K+\frac{1}{2}}}=C\,\frac{1}{\sqrt{m}}\,\sqrt{\frac{N}{K}}\,\exp\!\bigl(N\log N-m\log m-K\log K\bigr).$
Write $t:=p/m\geq 0$, so that $N=m(s+t)$ and K=m(s1+t)K=m(s-1+t). The exponent simplifies to
$N\log N-m\log m-K\log K=m\Bigl[(s+t)\log(s+t)-(s-1+t)\log(s-1+t)\Bigr].$
Define
$\Phi(t):=\log\Bigl(\frac{s^{s}}{(s-1)^{s-1}}\Bigr)+t\log\Bigl(\frac{s}{s-1}\Bigr)-\Bigl[(s+t)\log(s+t)-(s-1+t)\log(s-1+t)\Bigr].$
A direct computation gives $\Phi(0)=\Phi^{\prime}(0)=0$ and
$\Phi^{\prime\prime}(t)=\frac{1}{(s+t)(s-1+t)}>0,$
so $\Phi$ is convex with a global minimum at $t=0$. Hence $\Phi(t)\geq 0$ for all $t\geq 0$, i.e.
$(s+t)\log(s+t)-(s-1+t)\log(s-1+t)\leq\log(\zeta_{c}^{-1})+t\log(M).$
Substituting back yields
$\exp\bigl(N\log N-m\log m-K\log K\bigr)\leq\zeta_{c}^{-m}\,M^{p}.$
Also $\sqrt{N/K}\leq\sqrt{s/(s-1)}=\sqrt{M}$ and $\frac{p}{sm+p}\leq\frac{p}{sm}$. Combining these bounds gives
$R_{s,p}(m)\leq\frac{p}{sm}\cdot C\,\frac{1}{\sqrt{m}}\cdot\sqrt{M}\cdot\zeta_{c}^{-m}M^{p}\leq C_{s}\,p\,M^{p}\,\zeta_{c}^{-m}\,m^{-3/2},$
for a constant $C_{s}$ depending only on $s$. The extension to $m=0$ follows by enlarging $C_{s}$, since $R_{s,p}(0)=1$ and $(1+m)^{-3/2}=1$ at $m=0$. ∎
Lemma A.3 (Uniform one-term expansion in the tail region).
Fix $s\geq 2$ and $0<\theta<1$. There exists a constant $C_{s,\theta}>0$ such that for all integers $p\geq 1$ and $m\geq 1$ satisfying $p\leq\theta m$,
(A.4) $R_{s,p}(m)=A_{s,p}\,\zeta_{c}^{-m}\,m^{-3/2}\Bigl(1+\varepsilon_{s,p}(m)\Bigr),\qquad|\varepsilon_{s,p}(m)|\leq\frac{C_{s,\theta}}{m},$
where $A_{s,p}$ is as in (A.1).
Proof.
Write $t:=p/m$, so $0\leq t\leq\theta$, and write $p=tm$. Using the Gamma-form expression from the proof of Proposition A.1, consider
$\log\!\left(\frac{R_{s,p}(m)}{A_{s,p}\,\zeta_{c}^{-m}\,m^{-3/2}}\right).$
Applying the logarithmic Stirling expansion
$\log\Gamma(z)=\Bigl(z-\frac{1}{2}\Bigr)\log z-z+\frac{1}{2}\log(2\pi)+O(z^{-1})$
to the three Gamma factors with arguments $m(s+t)$, $m+1$, and m(s1+t)+1m(s-1+t)+1, one obtains
$\log R_{s,p}(m)=-m\log\zeta_{c}+\log A_{s,p}-\frac{3}{2}\log m+O(m^{-1}),$
uniformly for $t\in[0,\theta]$. Equivalently,
$\log\!\left(\frac{R_{s,p}(m)}{A_{s,p}\,\zeta_{c}^{-m}\,m^{-3/2}}\right)=O(m^{-1}),$
uniformly for $t\in[0,\theta]$. Here the terms of order $m$, the $\log m$ contribution, and the explicit $t$-dependent prefactor encoded in $A_{s,p}$ cancel identically; the remainder is uniform because all auxiliary functions of $t$ are smooth on the compact interval $0,\theta]$. Exponentiating yields
$\frac{R_{s,p}(m)}{A_{s,p}\,\zeta_{c}^{-m}\,m^{-3/2}}=1+O(m^{-1}),$
uniformly for $p\leq\theta m$, which is exactly (A.4). ∎
Appendix B Rank–one tail extraction in the Gram blocks
This appendix records the coefficient–level structure behind the rank–one decomposition in Theorem 4.6. Throughout we fix a symmetry sector $q\in\{1,\dots,s\}$, write
$p_{j}:=q+js,\qquad j\in\mathbb{N}_{0},$
and use the Gram vectors $\bm{v}^{(p)}$ from (3.8). Recall the block matrix $G^{(q)}(\zeta)$ determined coefficientwise by the Gram representation in (4.1), and its weighted conjugate
$\widetilde{G}^{(q)}(\zeta)=\mathcal{W}^{-1}G^{(q)}(\zeta)\mathcal{W}^{-1},$
where $\mathcal{W}=\mathrm{diag}(w_{j})$ and $w_{j}$ is given by (4.2).
B.1. Entry formula
Lemma B.1 (Entries of the Gram block).
For $j_{1}\leq j_{2}$ (and $\Delta:=j_{2}-j_{1}\geq 0$) one has
(B.1) $G^{(q)}_{j_{1}j_{2}}(\zeta)=\sum_{m=0}^{\infty}\frac{(p_{j_{2}}+sm)^{2}}{\sqrt{p_{j_{1}}p_{j_{2}}}}\,R_{s,p_{j_{1}}}(m+\Delta)\,R_{s,p_{j_{2}}}(m)\,\zeta^{\,2m+\Delta}.$
Consequently,
(B.2) $\widetilde{G}^{(q)}_{j_{1}j_{2}}(\zeta)=\frac{1}{w_{j_{1}}w_{j_{2}}}\,G^{(q)}_{j_{1}j_{2}}(\zeta),\qquad j_{1},j_{2}\in\mathbb{N}_{0}.$
Proof.
By definition, $G^{(q)}_{j_{1}j_{2}}(\zeta)=\langle\bm{v}^{(p_{j_{1}})},\bm{v}^{(p_{j_{2}})}\rangle$ in $\ell^{2}(\mathbb{N})$, and $\bm{v}^{(p)}$ is supported on indices $p+sm$. For $j_{1}\leq j_{2}$, the supports overlap precisely at $p_{j_{2}}+sm=p_{j_{1}}+s(m+\Delta),\ m\geq 0.$ Inserting (3.8) at these indices yields (B.1). The conjugated entry relation (B.2) is immediate from (4.4). ∎
B.2. The logarithmic tail and the singular vector
Write
$\eta:=\frac{\zeta}{\zeta_{c}}\in(0,1),\qquad\delta:=1-\eta^{2}=1-\frac{\zeta^{2}}{\zeta_{c}^{2}},\qquad L(\zeta)=\sum_{m\geq 1}\frac{\eta^{2m}}{m}=\log\frac{1}{1-\zeta^{2}/\zeta_{c}^{2}}.$
By Remark 2.8 (or Appendix A), for each fixed $p\geq 1$,
(B.3) $R_{s,p}(m)=A_{s,p}\,\zeta_{c}^{-m}\,m^{-3/2}\bigl(1+O(m^{-1})\bigr),\qquad m\to\infty,$
with an explicit constant $A_{s,p}>0$. Define the intrinsic amplitudes and their weighted realizations by
(B.4) $d^{(q)}_{j}:=\frac{s\,A_{s,p_{j}}}{\sqrt{p_{j}}},\qquad\widetilde{d}^{(q)}_{j}:=\frac{d^{(q)}_{j}}{w_{j}},\qquad j\in\mathbb{N}_{0}.$
Then $\widetilde{\bm{d}}^{(q)}:=(\widetilde{d}^{(q)}_{j})_{j\geq 0}\in\ell^{2}(\mathbb{N}_{0})$, and it is the vector appearing in Theorem 4.6.
Lemma B.2 (Uniform exponential gap away from the critical slope).
Fix $s\geq 2$ and $\theta>0$. There exist constants $C_{s,\theta},c_{s,\theta}>0$ such that for all integers $p,m\geq 1$ satisfying $p\geq\theta m$,
$R_{s,p}(m)\leq C_{s,\theta}\,p\,M^{p}\,\zeta_{c}^{-m}\,m^{-3/2}\,e^{-c_{s,\theta}p}.$
Proof.
Write $t:=p/m\in[\theta,\infty)$. By the Stirling estimate used in the proof of Proposition A.2, one has
$R_{s,p}(m)\leq C_{s}\,p\,m^{-3/2}\exp\bigl(m\Psi_{s}(t)\bigr),$
where
$\Psi_{s}(t):=(s+t)\log(s+t)-(s-1+t)\log(s-1+t).$
Equivalently,
$R_{s,p}(m)\leq C_{s}\,p\,M^{p}\,\zeta_{c}^{-m}\,m^{-3/2}\exp\Bigl(p\,\Xi_{s}(t)\Bigr),$
with
$\Xi_{s}(t):=\frac{\Psi_{s}(t)-\log(\zeta_{c}^{-1})}{t}-\log M.$
The function $\Xi_{s}$ is continuous on $\theta,\infty)$. The convexity argument used in Proposition A.2 yields
$\Psi_{s}(t)-t\log M\leq\log(\zeta_{c}^{-1}),$
with equality only at $t=0$, so $\Xi_{s}(t)<0$ for every $t>0$. Moreover, $\Xi_{s}(t)\to-\log M<0$ as $t\to\infty$. Hence
$\sup_{t\geq\theta}\Xi_{s}(t)=-c_{s,\theta}<0,$
which gives the stated bound. ∎
Lemma B.3 (Logarithmic tail with Hilbert–Schmidt remainder).
Let $\eta=\zeta/\zeta_{c}\in(0,1)$ and $L(\zeta)=\log\frac{1}{1-\eta^{2}}$. Define the rank–one Toeplitz kernel
$(K_{\eta})_{j_{1}j_{2}}:=\eta^{|j_{1}-j_{2}|}\,\widetilde{d}^{(q)}_{j_{1}}\,\widetilde{d}^{(q)}_{j_{2}},\qquad j_{1},j_{2}\in\mathbb{N}_{0}.$
Then, for every $0<\zeta<\zeta_{c}$, the operator
$R^{(q)}(\zeta):=\widetilde{G}^{(q)}(\zeta)-L(\zeta)\,K_{\eta}$
is Hilbert–Schmidt on $\ell^{2}(\mathbb{N}_{0})$, with $\sup_{\zeta<\zeta_{c}}\|R^{(q)}(\zeta)\|_{\mathrm{HS}}<\infty$. Moreover, as $\zeta\uparrow\zeta_{c}$, $R^{(q)}(\zeta)$ converges in Hilbert–Schmidt norm to a limit Hilbert–Schmidt operator $R_{*}^{(q)}$, hence also in operator norm.
In particular, for each fixed $(j_{1},j_{2})$ one has the entrywise expansion
(B.5) $\widetilde{G}^{(q)}_{j_{1}j_{2}}(\zeta)=\widetilde{d}^{(q)}_{j_{1}}\,\widetilde{d}^{(q)}_{j_{2}}\,\eta^{|j_{1}-j_{2}|}\,L(\zeta)+(R_{*}^{(q)})_{j_{1}j_{2}}+o(1),\qquad\zeta\uparrow\zeta_{c},$
and the error is $o(1)$ in $\ell^{2}$–matrix (Hilbert–Schmidt) sense.
Proof.
By symmetry of the Gram kernel, it is enough to treat the case $j_{1}\leq j_{2}$. Write $\Delta:=j_{2}-j_{1}\geq 0$. By Lemma B.1,
$\widetilde{G}^{(q)}_{j_{1}j_{2}}(\zeta)=\sum_{m\geq 0}U_{m}(j_{1},j_{2};\zeta),$
where
$U_{m}(j_{1},j_{2};\zeta):=\frac{1}{w_{j_{1}}w_{j_{2}}}\,\frac{(p_{j_{2}}+sm)^{2}}{\sqrt{p_{j_{1}}p_{j_{2}}}}\,R_{s,p_{j_{1}}}(m+\Delta)\,R_{s,p_{j_{2}}}(m)\,\zeta^{\,2m+\Delta}.$
Since $p_{j_{2}}\geq p_{j_{1}}$ and $p_{j_{2}}\geq\Delta$, the cutoff
$M_{j_{1}j_{2}}:=\max\{2p_{j_{1}},\,2p_{j_{2}},\,2\Delta,\,1\}$
reduces here to $M_{j_{1}j_{2}}=2p_{j_{2}}$. We split the sum into the initial range $0\leq m<2p_{j_{2}}$ and the tail $m\geq 2p_{j_{2}}$.
Tail region. If $m\geq 2p_{j_{2}}$, then $p_{j_{2}}\leq m/2$ and $p_{j_{1}}\leq(m+\Delta)/2$, so Lemma A.3 applies to both Raney factors (with $\theta=\frac{1}{2}$). More explicitly,
$R_{s,p_{j_{1}}}(m+\Delta)=A_{s,p_{j_{1}}}\zeta_{c}^{-m-\Delta}(m+\Delta)^{-3/2}\Bigl(1+O((m+1)^{-1})\Bigr),$
and
$R_{s,p_{j_{2}}}(m)=A_{s,p_{j_{2}}}\zeta_{c}^{-m}m^{-3/2}\Bigl(1+O((m+1)^{-1})\Bigr),$
uniformly for $m\geq 2p_{j_{2}}$. Since $j_{1}\leq j_{2}$, the quantities $p_{j_{1}},p_{j_{2}},\Delta$ are fixed along the tail sum, while
$(p_{j_{2}}+sm)^{2}=s^{2}m^{2}\Bigl(1+O((m+1)^{-1})\Bigr)$
and
$(m+\Delta)^{-3/2}m^{-3/2}=m^{-3}\Bigl(1+O((m+1)^{-1})\Bigr).$
Hence
$\frac{(p_{j_{2}}+sm)^{2}}{\sqrt{p_{j_{1}}p_{j_{2}}}}\,R_{s,p_{j_{1}}}(m+\Delta)\,R_{s,p_{j_{2}}}(m)\,\zeta^{2m+\Delta}=\frac{s^{2}A_{s,p_{j_{1}}}A_{s,p_{j_{2}}}}{\sqrt{p_{j_{1}}p_{j_{2}}}}\,\eta^{\Delta}\frac{\eta^{2m}}{m}\Bigl(1+O((m+1)^{-1})\Bigr).$
Using (B.4) and $m^{-1}=(m+1)^{-1}+O((m+1)^{-2})$, we obtain
(B.6) $U_{m}(j_{1},j_{2};\zeta)=\widetilde{d}^{(q)}_{j_{1}}\widetilde{d}^{(q)}_{j_{2}}\,\eta^{\Delta}\,\frac{\eta^{2m}}{m+1}+E_{m}^{\mathrm{tail}}(j_{1},j_{2};\zeta),$
with
(B.7) $|E_{m}^{\mathrm{tail}}(j_{1},j_{2};\zeta)|\leq C\,\widetilde{d}^{(q)}_{j_{1}}\widetilde{d}^{(q)}_{j_{2}}\,\eta^{\Delta}\,\frac{\eta^{2m}}{(m+1)^{2}},\qquad m\geq 2p_{j_{2}},$
where $C$ depends only on $s$ and $\beta$. Summing over $m\geq 2p_{j_{2}}$ shows that
$E^{\mathrm{tail}}_{j_{1}j_{2}}(\zeta):=\sum_{m\geq 2p_{j_{2}}}E_{m}^{\mathrm{tail}}(j_{1},j_{2};\zeta)$
defines a uniformly Hilbert–Schmidt kernel, since
$|E^{\mathrm{tail}}_{j_{1}j_{2}}(\zeta)|\leq C\,\widetilde{d}^{(q)}_{j_{1}}\widetilde{d}^{(q)}_{j_{2}}\sum_{m\geq 0}\frac{1}{(m+1)^{2}}\leq C^{\prime}\,\widetilde{d}^{(q)}_{j_{1}}\widetilde{d}^{(q)}_{j_{2}}$,
and $\widetilde{\bm{d}}^{(q)}\in\ell^{2}(\mathbb{N}_{0})$.
Initial range. Write
$E^{\mathrm{init}}_{j_{1}j_{2}}(\zeta):=\sum_{0\leq m<2p_{j_{2}}}U_{m}(j_{1},j_{2};\zeta).$
The term $m=0$ is estimated by Proposition A.2:
$|U_{0}(j_{1},j_{2};\zeta)|\leq C\,M^{-p_{j_{2}}}\,p_{j_{1}}^{-1-\beta}p_{j_{2}}^{1-\beta}(1+\Delta)^{-3/2}.$
For $1\leq m<2p_{j_{2}}$, Proposition A.2 applies to $R_{s,p_{j_{1}}}(m+\Delta)$, while Lemma B.2 with $\theta=\frac{1}{2}$ applies to $R_{s,p_{j_{2}}}(m)$ because $p_{j_{2}}\geq m/2$. Using $\zeta^{2m+\Delta}\leq\zeta_{c}^{2m+\Delta}$, one gets
$|U_{m}(j_{1},j_{2};\zeta)|\leq C\,e^{-cp_{j_{2}}}\,p_{j_{1}}^{-1-\beta}p_{j_{2}}^{-1-\beta}\,\frac{(p_{j_{2}}+sm)^{2}}{(m+1)^{3/2}(m+\Delta+1)^{3/2}}.$
Since $m<2p_{j_{2}}$, we have $p_{j_{2}}+sm\leq(1+2s)p_{j_{2}}$, and therefore
$\sum_{m=1}^{2p_{j_{2}}-1}|U_{m}(j_{1},j_{2};\zeta)|\leq C\,e^{-cp_{j_{2}}}\,p_{j_{1}}^{-1-\beta}p_{j_{2}}^{1-\beta}\sum_{m\geq 1}\frac{1}{(m+1)^{3}}\leq C^{\prime}\,e^{-cp_{j_{2}}}\,p_{j_{1}}^{-1-\beta}p_{j_{2}}^{1-\beta}.$
Hence
$|E^{\mathrm{init}}_{j_{1}j_{2}}(\zeta)|\leq C^{\prime\prime}\,e^{-c^{\prime}p_{j_{2}}}\,p_{j_{1}}^{-1-\beta}p_{j_{2}}^{1-\beta},$
for constants $C^{\prime\prime},c^{\prime}>0$ independent of $\zeta$. This bound is square summable in $(j_{1},j_{2})$, so $E^{\mathrm{init}}(\zeta)$ is uniformly Hilbert–Schmidt. Moreover, for each fixed $(j_{1},j_{2})$ the sum is finite and therefore converges termwise as $\zeta\uparrow\zeta_{c}$.
Summing (B.6) over $m\geq 2p_{j_{2}}$ gives
$\sum_{m\geq 2p_{j_{2}}}U_{m}(j_{1},j_{2};\zeta)=\widetilde{d}^{(q)}_{j_{1}}\widetilde{d}^{(q)}_{j_{2}}\,\eta^{\Delta}\sum_{m\geq 2p_{j_{2}}}\frac{\eta^{2m}}{m+1}+E^{\mathrm{tail}}_{j_{1}j_{2}}(\zeta).$
Reinstating the omitted harmonic terms produces the correction
$E^{\mathrm{harm}}_{j_{1}j_{2}}(\zeta):=\widetilde{d}^{(q)}_{j_{1}}\widetilde{d}^{(q)}_{j_{2}}\,\eta^{\Delta}\sum_{m=0}^{2p_{j_{2}}-1}\frac{\eta^{2m}}{m+1}.$
Since
$\sum_{m=0}^{2p_{j_{2}}-1}\frac{\eta^{2m}}{m+1}\leq 1+\log(1+2p_{j_{2}}),$
we obtain
$|E^{\mathrm{harm}}_{j_{1}j_{2}}(\zeta)|\leq C\,\widetilde{d}^{(q)}_{j_{1}}\widetilde{d}^{(q)}_{j_{2}}\bigl(1+\log(1+p_{j_{2}})\bigr).$
Because $\widetilde{d}^{(q)}_{j}\asymp p_{j}^{-1-\beta}$, the kernel on the right-hand side is Hilbert–Schmidt:
$\sum_{j_{1},j_{2}\geq 0}\widetilde{d}^{(q)}_{j_{1}}{}^{2}\widetilde{d}^{(q)}_{j_{2}}{}^{2}\bigl(1+\log(1+p_{j_{2}})\bigr)^{2}<\infty.$
Thus
$\widetilde{G}^{(q)}_{j_{1}j_{2}}(\zeta)=\widetilde{d}^{(q)}_{j_{1}}\widetilde{d}^{(q)}_{j_{2}}\,\eta^{\Delta}\sum_{m\geq 0}\frac{\eta^{2m}}{m+1}+E^{\mathrm{init}}_{j_{1}j_{2}}(\zeta)+E^{\mathrm{tail}}_{j_{1}j_{2}}(\zeta)+E^{\mathrm{harm}}_{j_{1}j_{2}}(\zeta).$
Since $\sum_{m\geq 0}\frac{\eta^{2m}}{m+1}=\eta^{-2}L(\zeta)$, we write
$\eta^{-2}L(\zeta)=L(\zeta)+\bigl(\eta^{-2}-1\bigr)L(\zeta).$
The coefficient $\bigl(\eta^{-2}-1\bigr)L(\zeta)$ stays bounded as $\eta\uparrow 1$, and $K_{\eta}$ is uniformly Hilbert–Schmidt. Therefore the extra term $\bigl(\eta^{-2}-1\bigr)L(\zeta)\,K_{\eta}$ is uniformly Hilbert–Schmidt as well. Collecting all Hilbert–Schmidt pieces into $R^{(q)}(\zeta)$ yields
$\widetilde{G}^{(q)}(\zeta)=L(\zeta)\,K_{\eta}+R^{(q)}(\zeta),$
with $\sup_{\zeta<\zeta_{c}}\|R^{(q)}(\zeta)\|_{\mathrm{HS}}<\infty$.
Finally, each constituent of $R^{(q)}(\zeta)$ has a pointwise limit as $\zeta\uparrow\zeta_{c}$, and the dominating Hilbert–Schmidt kernels displayed above are independent of $\zeta$. Dominated convergence therefore yields
$R^{(q)}(\zeta)\to R_{*}^{(q)}\qquad\text{in Hilbert--Schmidt norm as }\zeta\uparrow\zeta_{c},$
hence also in operator norm. ∎
B.3. From entrywise asymptotics to a rank–one operator
The prefactor $\eta^{|j_{1}-j_{2}|}$ in (B.5) tends to $1$ as $\zeta\uparrow\zeta_{c}$. The next lemma quantifies that, after weighting, it can be absorbed into the bounded remainder.
Lemma B.4 (Removal of the Toeplitz prefactor).
Let $D=(\widetilde{d}^{(q)}_{j})_{j\geq 0}\in\ell^{2}(\mathbb{N}_{0})$ and define operators
$K_{\eta}:\ (K_{\eta})_{j_{1}j_{2}}:=\eta^{|j_{1}-j_{2}|}D_{j_{1}}D_{j_{2}},\qquad K_{1}:=D\otimes D^{*},\qquad 0<\eta\leq 1.$
Then $K_{\eta}-K_{1}\to 0$ in Hilbert–Schmidt norm as $\eta\uparrow 1$. In particular,
$\|K_{\eta}-K_{1}\|\to 0,\qquad L(\zeta)\,\|K_{\eta}-K_{1}\|\to 0\quad\text{as }\zeta\uparrow\zeta_{c}.$
Proof.
By (B.4), (4.2) and the explicit constant $A_{s,p}=\frac{p\,M^{p}}{\sqrt{2\pi s(s-1)}}$ in (A.1), one has
$D_{j}=\widetilde{d}^{(q)}_{j}=c_{s}\,p_{j}^{-1-\beta},\qquad c_{s}:=\sqrt{\frac{s}{2\pi(s-1)}},$
hence $a_{j}:=D_{j}^{2}\leq C(1+j)^{-2-2\beta}$ and $\sum_{j\geq 0}a_{j}<\infty$.
Set $b_{d}:=\sum_{j\geq 0}a_{j}a_{j+d}$ for $d\geq 0$. By symmetry,
$\|K_{\eta}-K_{1}\|_{\mathrm{HS}}^{2}=\sum_{j_{1},j_{2}\geq 0}(1-\eta^{|j_{1}-j_{2}|})^{2}a_{j_{1}}a_{j_{2}}\asymp\sum_{d\geq 0}(1-\eta^{d})^{2}b_{d},$
up to an absolute factor. Moreover, since $a_{j+d}\leq C(1+d)^{-2-2\beta}$ for all $j\geq 0$,
$b_{d}=\sum_{j\geq 0}a_{j}a_{j+d}\leq C(1+d)^{-2-2\beta}\sum_{j\geq 0}a_{j}\leq C^{\prime}(1+d)^{-2-2\beta}.$
Fix $N\in\mathbb{N}$. Using $1-\eta^{d}\leq d(1-\eta)$ for $d\leq N$ and $1-\eta^{d}\leq 1$ for $d>N$, we get
$\sum_{d\geq 0}(1-\eta^{d})^{2}b_{d}\leq(1-\eta)^{2}\sum_{d=0}^{N}d^{2}b_{d}+\sum_{d>N}b_{d}\leq C^{\prime}(1-\eta)^{2}\sum_{d=1}^{N}d^{-2\beta}+C^{\prime}\sum_{d>N}d^{-2-2\beta}.$
The last tail is $O(N^{-1-2\beta})$. The partial sum satisfies $\sum_{d=1}^{N}d^{-2\beta}=O(N^{1-2\beta})$ if $\beta<\frac{1}{2}$, $O(\log N)$ if $\beta=\frac{1}{2}$, and $O(1)$ if $\beta>\frac{1}{2}$. Choose $N:=\lfloor(1-\eta)^{-1}\rfloor$ to balance the two terms. This yields
$\begin{cases}O\!\bigl((1-\eta)^{1+2\beta}\bigr),&0<\beta<\frac{1}{2},\\[2.0pt] O\!\bigl((1-\eta)^{2}\log\frac{1}{1-\eta}\bigr),&\beta=\frac{1}{2},\\[2.0pt] O\!\bigl((1-\eta)^{2}\bigr),&\beta>\frac{1}{2},\end{cases}\qquad\eta\uparrow 1.$
In particular,
$\|K_{\eta}-K_{1}\|_{\mathrm{HS}}=O\!\bigl((1-\eta)^{\min\{1,\,\frac{1}{2}+\beta\}}\bigr)$
up to an extra factor $\sqrt{\log\frac{1}{1-\eta}}$ when $\beta=\frac{1}{2}$. Since $\min\{1,\frac{1}{2}+\beta\}\geq\min\{1,\beta\}$ for all $\beta>0$, this is the required Hilbert–Schmidt estimate.
Finally, $\|K_{\eta}-K_{1}\|\leq\|K_{\eta}-K_{1}\|_{\mathrm{HS}}$ and $L(\zeta)\sim\log\frac{1}{1-\eta}$ with $(1-\eta)^{\gamma}\log\frac{1}{1-\eta}\to 0$ for any $\gamma>0$, yield $L(\zeta)\|K_{\eta}-K_{1}\|\to 0$. This proves the lemma. ∎
Appendix C Hypergeometric continuation of $\mathcal{G}_{p}$
We give a proof of the hypergeometric representation stated in Proposition 5.4. We keep the notation $\{\alpha_{i}\}_{i=1}^{2s}$ and $\{\beta_{j}\}_{j=1}^{2s-1}$ from (5.6).
Proof of Proposition 5.4.
Write
$\mathcal{G}_{p}(u)=\sum_{m\geq 0}a_{m}\,u^{m},\qquad a_{m}:=R_{s,p}(m)^{2},\qquad R_{s,p}(m)=\frac{p}{sm+p}\binom{sm+p}{m}$.
Using
$R_{s,p}(m)=p\,\frac{\Gamma(sm+p)}{\Gamma(m+1)\Gamma((s-1)m+p+1)},$
one computes
$\frac{R_{s,p}(m+1)}{R_{s,p}(m)}=\frac{\prod_{k=0}^{s-1}(sm+p+k)}{(m+1)\prod_{l=1}^{s-1}((s-1)m+p+l)}.$
Squaring and factorizing the linear terms gives
$\frac{a_{m+1}}{a_{m}}=\zeta_{c}^{-2}\,\frac{\prod_{k=0}^{s-1}\left(m+\frac{p+k}{s}\right)^{2}}{(m+1)^{2}\prod_{l=1}^{s-1}\left(m+\frac{p+l}{s-1}\right)^{2}},$
because
$\frac{s^{2s}}{(s-1)^{2s-2}}=\zeta_{c}^{-2}.$
Now consider the generalized hypergeometric series
$H(u):={}_{2s}F_{2s-1}\!\left(\begin{matrix}\alpha_{1},\ldots,\alpha_{2s}\\ \beta_{1},\ldots,\beta_{2s-1}\end{matrix}\Big|\,\zeta_{c}^{-2}u\right),$
with parameter multisets
$\{\alpha_{i}\}=2\times\left\{\frac{p+k}{s}:k=0,\ldots,s-1\right\},\qquad\{\beta_{j}\}=\{1\}\cup 2\times\left\{\frac{p+l}{s-1}:l=1,\ldots,s-1\right\}.$
Its coefficients $b_{m}$ satisfy $b_{0}=1$ and the standard ratio formula
$\frac{b_{m+1}}{b_{m}}=\zeta_{c}^{-2}\,\frac{\prod_{i=1}^{2s}(m+\alpha_{i})}{(m+1)\prod_{j=1}^{2s-1}(m+\beta_{j})}.$
By the chosen parameter sets, this is exactly the ratio displayed above for $a_{m+1}/a_{m}$. Since $a_{0}=R_{s,p}(0)^{2}=1=b_{0}$, the sequences coincide for all $m\geq 0$, and therefore $\mathcal{G}_{p}(u)=H(u)$. This proves (5.5). ∎
Appendix D The logarithmic coefficient at the branch point
This appendix computes the coefficient $B(\zeta_{c}^{2})$ in the resonant local expansion (5.7) of $\mathcal{G}_{p}$ at the branch point $u=\zeta_{c}^{2}$. Recall that the radius of convergence of $\mathcal{G}_{p}$ equals $\zeta_{c}^{2}$.
Proposition D.1 (Explicit value of $B(\zeta_{c}^{2})$).
For all integers $s\geq 2$ and $p\geq 1$, the coefficient $B(\zeta_{c}^{2})$ in (5.7) satisfies
(D.1) $B(\zeta_{c}^{2})=-\frac{p^{2}}{4\pi}\,\frac{s^{\,2p-1}}{(s-1)^{\,2p+1}}.$
In particular, $B(\zeta_{c}^{2})<0$.
Proof.
Write $\mathcal{G}_{p}(u)=\sum_{m\geq 0}R_{s,p}(m)^{2}\,u^{m}$ and introduce the scaled variable $\xi:=u/\zeta_{c}^{2}$, so that
$\mathcal{G}_{p}(\zeta_{c}^{2}\xi)=\sum_{m\geq 0}a_{m}\xi^{m},\qquad a_{m}:=R_{s,p}(m)^{2}\,\zeta_{c}^{2m}.$
By Proposition A.1, one has the refined asymptotic
(D.2) $a_{m}=C_{s,p}\,m^{-3}+O(m^{-4}),\qquad C_{s,p}:=\frac{p^{2}}{2\pi}\,\frac{s^{\,2p-1}}{(s-1)^{\,2p+1}}.$
On the other hand, the polylogarithm $\mathrm{Li}_{3}(\xi)=\sum_{m\geq 1}m^{-3}\xi^{m}$ has the classical singular expansion at $\xi=1$ (see, e.g., DLMF §25.12 [18])
$\mathrm{Li}_{3}(\xi)=\text{(analytic at $\xi=1$)}-\tfrac{1}{2}(1-\xi)^{2}\log(1-\xi)+O\!\bigl((1-\xi)^{2}\bigr),\qquad\xi\to 1.$
Define the remainder series
$H(\xi):=\mathcal{G}_{p}(\zeta_{c}^{2}\xi)-C_{s,p}\,\mathrm{Li}_{3}(\xi)=\sum_{m\geq 1}\Bigl(a_{m}-C_{s,p}m^{-3}\Bigr)\xi^{m}.$
By (D.2), the coefficients satisfy $a_{m}-C_{s,p}m^{-3}=O(m^{-4})$, hence $\sum_{m\geq 1}m^{2}\bigl|a_{m}-C_{s,p}m^{-3}\bigr|<\infty$. Therefore the series for $H(\xi)$, $H^{\prime}(\xi)$ and $H^{\prime\prime}(\xi)$ converge absolutely at $\xi=1$, so $H$ extends to a $C^{2}$-function on $[0,1]$ and therefore
$H(\xi)=H(1)+H^{\prime}(1)(\xi-1)+\tfrac{1}{2}H^{\prime\prime}(1)(\xi-1)^{2}+o\bigl((1-\xi)^{2}\bigr),\qquad\xi\to 1.$
In particular, $H$ contributes no $(1-\xi)^{2}\log(1-\xi)$ term.
Consequently, the $(1-\xi)^{2}\log(1-\xi)$ coefficient in $\mathcal{G}_{p}(\zeta_{c}^{2}\xi)$ is exactly $\tfrac{1}{2}C_{s,p}$, coming from $C_{s,p}\mathrm{Li}_{3}(\xi)$. Comparing with (5.7) yields $B(\zeta_{c}^{2})=-\tfrac{1}{2}C_{s,p}$, which is (D.1). ∎
[…]
For all integers $s\geq 2$ and $p\geq 1$ one has $\pi\,B(\zeta_{c}^{2})\in\mathbb{Q}$ and $\pi\,B(\zeta_{c}^{2})<0$. Consequently, the edge value in Theorem 5.13 satisfies $\rho_{p}(\zeta_{c}^{2})\in\pi^{-1}\mathbb{Q}$.